### v1

# A Dimca-Greuel type inequality for foliations

Abstract.
Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$. We prove the following Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}[[x,y]]$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$. As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel’s conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
Key words and phrases:
Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
2020 Mathematics Subject Classification:
Primary 32S65 - 32M25
Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done (Spanish grant PID2019-105896GB-I00 funded). This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979.
1. Introduction
Let $\mathbb{C}\{x,y\}$ be the ring of complex convergent power series in two variables. Consider the germ of isolated plane curve singularity $C$ defined by the non-unit $f(x,y)\in\mathbb{C}\{x,y\}$. The Milnor number of $C$ is
$\mu(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(\partial_{x}f,\partial_{y}f)$
and its Tjurina number is
$\tau(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(f,\partial_{x}f,\partial_{y}f).$
Is clear from their definitions that $\tau(C)\leq\mu(C)$.
In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity. The question was affirmatively answered in [1], [10], and [15] for the irreducible case and in [2] for the reduced case.
It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs. More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$. Moreover if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9], [11]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$. If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
(1) $\frac{\mu_{p}(\mathcal{F})}{\tau_{p}(\mathcal{F},\mathcal{B}_{0})}$
equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$.
If the foliation $\mathcal{F}$ is of the second type and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied). It is natural to wonder if we can improve the bound $2$ for the ratio (1).
In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, given by the $1$-form
$\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$ then we get a bound of the Dimca-Greuel type for foliations.
Our main theorem is
Theorem A.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
$\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
We will prove Theorem A in Section 3.
As a consequence of Theorem A we deduce (see Corollary 3.5) the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [15].
2. Multiplicity of a foliation along a divisor of separatrices
Throughout this text $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$. Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
(2) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, page 159]. If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
(3) $\theta(t)=\begin{cases}-\frac{Q(\gamma(t))}{x^{\prime}(t)}&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \frac{P(\gamma(t))}{y^{\prime}(t)}&\text{if $y(t)\neq 0$}.\end{cases}$
We extend the notion of multiplicity of $\mathcal{F}$ along any nonempty divisor $\mathcal{B}=\sum_{B}a_{B}B$ of separatrices of $\mathcal{F}$ at $p$ in the following way:
(4) $\mu_{p}(\mathcal{F},\mathcal{B})=\left(\displaystyle\sum_{B}a_{B}\mu_{p}(\mathcal{F},B)\right)-\deg(\mathcal{B})+1.$
By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$. In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p. 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$). Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
(5) $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
Since $\mu(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we get
(6) $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
Denote by $GSV_{p}(\mathcal{F},\mathcal{B})$ the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to the effective divisor $\mathcal{B}$ of separatrices of $\mathcal{F}$. See [12] for details.
Remark 2.1.
Let $B:f(x,y)=0$ be a separatrix of $\mathcal{F}$ and let $\mathcal{G}_{f}$ be the hamiltonian foliation defined by $df=0$. From (2) and (3), we have
(7) $\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$. By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$. But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$
3. Proof of Theorem A
Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
(8) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (8):
(9) $\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$). Denote by $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$. If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([15]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$. By [15, Claim 4.4] we have
(10) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},$
for $\nu_{p}(C)\geq 2$.
We will need the following lemma.
Lemma 3.1.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$, then
(11) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Moreover, the equality holds if and only if $B$ is smooth.
Proof.
Suppose first $\nu_{p}(B)\geq 2$. By [5, Equality (1.2), p. 291] we get
(12) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
where
(13) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
On the other hand by the behaviour under blow-up of the GSV index [3, p.30] we get
(14) $GSV_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
and by [7, Proposition 6.2] we have
(15) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
After [15, Equation (4)] we get:
(16) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
where $\alpha$ was defined in (10). From (15), (14), and (16) we deduce
(17) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}+\alpha.$
By (12), (14) and (17) it holds
(18) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$ $=$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
-- $\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha.$
Hence, if $B$ is a singular curve then the lemma follows by (10).
Suppose now that $B$ is a non-singular curve. By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$. Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$. Hence $Q=xh$ for some formal power series $h(x,y)\in\mathbb{C}[[x,y]]$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$. Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
(19) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,x)=\mu_{p}(\mathcal{F},B).$
Moreover by (15) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$. Thus if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality.
Moreover, if (11) is an equality, then $B$ is necessarily smooth. Indeed if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (18) we have that (11) is a strict inequality that is a contradiction. Hence the lemma follows. ∎
Corollary 3.2.
Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$ after a composition of $\ell$ blows-up then
(20) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
Proof.
It follows applying Lemma 3.1 $\ell$ times. ∎
Remark 3.3.
By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
(21) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
In addition, by (19) if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
Remark 3.4.
If $\mathcal{F}$ is a non-dicritical second type foliation at $p\in\mathbb{C}^{2}$, and $B$ is the only separatrix of $\mathcal{F}$, then $\mathcal{F}$ is a generalized curve foliation. In fact, assume by contradiction that $\mathcal{F}$ admits a saddle-node singularity in its reduction process with strong separatrix on the exceptional divisor. Without lost of generality, we can suppose that $\mathcal{F}$ is a reduced foliation after a blow-up $\sigma$ at $p$. Since $\mathcal{F}$ is of second type, after a blow-up, the strong separatrix of the strict transform $\tilde{\mathcal{F}}$ of $\mathcal{F}$ through $q\in\sigma^{-1}(p)$ is contained in $E=\sigma^{-1}(p)$, thus the Camacho-Sad index is $CS_{q}(\tilde{\mathcal{F}},E)=0$, but it is a contradiction with Camacho-Sad’s formula
$CS_{q}(\tilde{\mathcal{F}},E)=E\cdot E=-1.$
Proof of Theorem A: If $B$ is a non-singular curve then, by Remark 3.3, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$ and the theorem follows.
Let us suppose now that $B$ is a singular separatrix, that is $\nu_{p}(B)\geq 2$. After Corollary 3.2 we can suppose, without lost of generality that the number of blows-up $n$ needed in the process of the reduction of the singularity of $B$ is one. In particular, after a blow-up $\sigma$ the strict transform $\tilde{B}$ of $B$ is smooth, that is the algebraic multiplicity of $\tilde{B}$ is one.
Applying Lemma 3.1 and after (14) we get
(22) $[3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})]-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
Since $\tilde{B}$ is smooth, from (22) and the equality $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{q}(\tilde{\mathcal{F}},\tilde{B})$ (see Remark 3.3) we obtain
(23) $-\tau_{q}(\tilde{\mathcal{F}},\tilde{B})-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
Moreover, by the definition of $m_{p}(\mathcal{F})$ and after [11, Proposition 3.3], if $\mathcal{B}=B+\mathcal{B}^{\prime}$ is a balanced divisor of separatrices of $\mathcal{F}$ (for some divisor $\mathcal{B}^{\prime}$, if $B$ is the only separatrix of $\mathcal{F}$ we consider for $\mathcal{B}^{\prime}$ the empty divisor) then we have
$m_{p}(\mathcal{F})-\nu_{p}(B)=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{B}^{\prime})-1+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical}\\ \nu_{p}(\mathcal{B}^{\prime})+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is dicritical},\end{array}\right.$
where $\xi_{p}(\mathcal{F})$ is the tangency excess of $\mathcal{F}$ (see for example [7, Definition 2.3]). So $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$ except in the case that $\mathcal{F}$ is a foliation of the second type ($\xi_{p}(\mathcal{F})=0$) and $B$ is its only separatrix ($\nu_{p}(\mathcal{B}^{\prime})=0$) in which case we get $m_{p}(\mathcal{F})-\nu_{p}(B)=-1$. In the case $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$, from (23) we obtain $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. Otherwise, assume that $\mathcal{F}$ is of the second type and that $B$ is its only separatrix. By Remark 3.4, $\mathcal{F}$ is a generalized curve foliation, so $GSV_{p}(\mathcal{F},B)=0$. Moreover, since $\tilde{B}$ is smooth then the left-term of (11) is zero, so the theorem follows by Lemma (3.1) (since $B$ is singular the inequality (11) is strict) and Remark 3.3 ($\tilde{B}$ is smooth). $\Box$
From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [15]:
Corollary 3.5.
Let $f(x,y)\in\mathbb{C}[[x,y]]$ an irreducible power series. Put $C:f(x,y)=0$ the branch defined by $f$. Then
$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
Proof.
It is a consequence of Remark 2.1 and of applying Theorem A to foliation $\mathcal{F}:df=0$. ∎
Remark 3.6.
In [15] Wang proved that the number $3\mu_{p}(B)-4\tau_{p}(B)$ is a non-negative number, non-decreasing under blowups and strictly increasing unless the curve $B$ is non-singular. The result of Wang is an special case of Lemma 3.1, taking $\mathcal{F}=df$ where $B$ is given by $f(x,y)=0$. Lemma 3.1 and Theorem A generalizes this for the number $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$.
Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
Corollary 3.7.
Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ be a balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
(24) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
Proof.
By (6) we have $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}$. Hence we will prove that
(25) $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
Suppose that $\mathcal{B}_{0}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
$\sum_{j=1}^{\ell}\left(3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})\right)\leq-\ell<-\ell+1.$
Thus
$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})<-\ell+1=3\ell-4\ell+(4-3)$
and
$3\left(\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-\ell+1\right)<4\left(\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})-\ell+1\right),$
so
$\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
The following example illustrates Corollary 3.7.
Example 3.8.
Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$.
Let $C=\{x=0\}$, $\mathcal{B}_{0}=C+C_{0,1}+C_{1,1}$, $\mathcal{B}_{\infty}=C_{2,1}$, then $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices for $\mathcal{F}$. We have
$\mu_{0}(\mathcal{F},\mathcal{B}_{0})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{0,1})+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+29+21-2+1=54,$
$\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})=\mu_{0}(\mathcal{F},C_{2,1})=21,$
and
$T_{0}(\mathcal{F},\mathcal{B}_{0})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{0,1})+\tau_{0}(\mathcal{F},C_{1,1})=5+28+21=54.$
Hence
$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},\mathcal{B}_{0})-\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})+1=54-21+1=34,$
and
$\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B}_{0})}=\frac{34}{54}<\frac{4}{3}.$
We conclude the paper with the following question:
Question: Is the inequality (24) true if instead of $T_{p}(\mathcal{F},\mathcal{B}_{0})$ we write $\tau_{p}(\mathcal{F},\mathcal{B}_{0})$?

### v2

@@
 Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$.
-We prove the following Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}[[x,y]]$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$.
-As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel’s conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
+Under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we prove the Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$.
+As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
 Key words and phrases:
@@
 Is clear from their definitions that $\tau(C)\leq\mu(C)$.
-In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity.
+In [7, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity.
 The question was affirmatively answered in [1], [11], and [19] for the irreducible case and in [2] for the reduced case.
@@
 More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$.
-Moreover if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9], [11]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$.
+The algebraic multiplicity $\nu_{p}(\mathcal{F})$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
+Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [10], [12]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$.
 If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
@@
 equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$.
-If the foliation $\mathcal{F}$ is of the second type and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied).
+If the foliation $\mathcal{F}$ is of the second type (see [10, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [9, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [15] where the general case of hypersurfaces is studied).
 It is natural to wonder if we can improve the bound $2$ for the ratio (1).
-In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, given by the $1$-form
+In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [8, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
 $\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
 then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
-However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$ then we get a bound of the Dimca-Greuel type for foliations.
+However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$, then, under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we will obtain a bound of the Dimca-Greuel type for foliations.
+More specifically, in Foliation theory, two type of separatrices of $\mathcal{F}$ are distinguished: isolated separatrices $\text{ISO}_{p}(\mathcal{F})$ and dicritical separatrices $\text{Dic}_{p}(\mathcal{F})$.
+Isolated separatrices of $\mathcal{F}$ correspond to separatrices that are images, by the reduction process of singularities $\pi$ of $\mathcal{F}$, of separatrices of non-degenerate singularities, or weak or strong separatrices of saddle-nodes singularities in a reduction of singularities of $\mathcal{F}$.
+On the other hand, dicritical separatrices correspond to separatrices that are images, by $\pi$, of separatrices that intersect transversally with a dicritical divisor in the last step of the reduction of singularities of $\mathcal{F}$.
+For more details about these definitions, see Section 3.
+In order to establish an inequality of Dimca-Greuel type, it is necessary to introduce a number $\text{DG}_{p}(\mathcal{F},B)$ that depends on the reduction of singularities of the pair $(\mathcal{F},B)$ and on the type of separatrix that $B$ is, see Definition 21.
+We will see that the negativity of this number is crucial in order to obtain an inequality of the Dimca-Greuel type for foliations.
 Our main theorem is
 Theorem A.
-Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
-$\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
-We will prove Theorem A in Section 3.
+Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\text{DG}_{p}(\mathcal{F},B)<0$, then
+(2) $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
+We will prove Theorem A in Section 4. In Example 4.3, we will show that we cannot removed the hypothesis over $\text{DG}_{p}(\mathcal{F},B)$.
 As a consequence of Theorem A we deduce (see Corollary 5.1) the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
 2. Multiplicity of a foliation along a divisor of separatrices
-Throughout this text $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
-$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
-The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
+Throughout this note $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
+(3) $v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
 Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
-Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$.
+We denote by $\text{Sep}_{p}(\mathcal{F})$ the set of all separatrices of $\mathcal{F}$ at $p$.
+Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by a vector field $v$ as in (3) and $B\in\text{Sep}_{p}(\mathcal{F})$.
 Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
 (4) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
-where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, page 159].
+where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [5, p.
+159].
 If $\omega=P(x,y)dx+Q(x,y)dy$ is a 1-form inducing $\mathcal{F}$ and $\gamma(t)=(x(t),y(t))$, then
@@
 By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$.
-In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p.
+In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [14, Equation (2.2), p.
 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
@@
 (7) $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
-Since $\mu(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we get
+Since $\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we have
 (8) $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
@@
 where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$.
-By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$.
-But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$
+By [8, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$.
+But, since $B$ is irreducible, after [8, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [8, Proposition 6.2], we obtain
+(10) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
+Consequently when $B$ is quasihomogeneous by ([17], Satz p.123 a) and d)] and (10) or by [20, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
+3. Reduction of singularities of foliations and their separatrices
+We say that $p\in\mathbb{C}^{2}$ is a reduced singularity for $\mathcal{F}$ if the linear part $\text{D}v(p)$ of the vector field $v$ in (3) is non-zero and has eigenvalues $\lambda_{1},\lambda_{2}\in\mathbb{C}$ fitting in one of the cases:
+  1. (i)
+    $\lambda_{1}\lambda_{2}\neq 0$ and $\lambda_{1}\lambda_{2}\not\in\mathbb{Q}^{+}$ (non-degenerate);
+  2. (ii)
+    $\lambda_{1}\neq 0$ and $\lambda_{2}\neq 0$ (saddle-node singularity).
+In the case $(i)$, there is a system of coordinates $(x,y)$ in which $\mathcal{F}$ is defined by the equation
+(11) $\omega=x(\lambda_{1}+a(x,y))dy-y(\lambda_{2}+b(x,y))dx,$
+where $a(x,y),b(x,y)\in{\mathbb{C}}[[x,y]]$ are non-units, so that $\text{Sep}_{p}(\mathcal{F})$ is formed by two transversal analytic branches given by $\{x=0\}$ and $\{y=0\}$.
+In the case $(ii)$, up to a formal change of coordinates, the saddle-node singularity is given by a 1-form of the type
+(12) $\omega=x^{k+1}dy-y(1+\lambda x^{k})dx,$
+where $\lambda\in\mathbb{C}$ and $k\in\mathbb{Z}^{+}$ are invariants after formal changes of coordinates (see [16, Proposition 4.3]).
+The curve $\{x=0\}$ is an analytic separatrix, called strong separatrix, whereas $\{y=0\}$ corresponds to a possibly formal separatrix, called weak separatrix.
+The integer $k+1>1$ is called tangency index of $\mathcal{F}$ with respect to the weak separatrix.
+Let $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ be a composition of blow-ups maps.
+The divisor $\mathcal{D}$ is a finite union of components which are embedding projective lines, crossing normally at corners.
+If $\mathcal{F}$ is the foliation given by the 1-form $\omega$, we denote ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ the strict transform of $\mathcal{F}$, the germ of foliation on $(\tilde{X},\mathcal{D})$ defined locally by $\pi^{*}(\omega)$, after cancelling one-dimensional singular components.
+For a uniform analysis, we include the possibility of $\pi$ being the identity map and, abusing notation, we set in this case $\tilde{X}=\mathbb{C}^{2}$, $\mathcal{D}=\{p\}$ and $\tilde{\mathcal{F}}=\mathcal{F}$.
+With respect to the divisor $\mathcal{D}$, the foliation $\tilde{\mathcal{F}}$ at a point $q\in\mathcal{D}$ can be:
+  • Regular, if there are local analytic coordinates $(x,y)$ at $q$ such that $\mathcal{D}\subset\{xy=0\}$ and ~$\tilde{\mathcal{F}}:dx=0$.
+  • Singular; if it is not regular.
+  • Reduced, if $q$ is a reduced singularity for $\tilde{\mathcal{F}}$ and $\mathcal{D}\subset\text{Sep}_{q}(\tilde{\mathcal{F}})$.
+When $\mathcal{D}=\{p\}$, these notions coincide with the ordinary concepts of regular point, singular point and reduced singularity.
+We say that $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ is a reduction of singularities for $\mathcal{F}$ if all points $q\in\mathcal{D}$ are either regular or reduced singularities.
+There always exists a reduction of singularities [18].
+Besides, there exists a minimal one, in the sense that it factorizes any other reduction of singularities by an additional sequence of blow-ups.
+All along this note, reductions of singularities are supposed to be minimal.
+Given a germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ with a fixed reduction process $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},0)$ for $\mathcal{F}$, a component $D\subset\mathcal{D}$ can be:
+  • Non-dicritical, if $D$ is $\tilde{\mathcal{F}}$-invariant.
+In this case $D$ contains a finite number of reduced singularities.
+Each non-corner singularity carries a separatrix transversal to $D$, whose projection by $\pi$ is a curve in $\text{Sep}_{p}(\mathcal{F})$.
+  • Dicritical, if $D$ is not $\tilde{\mathcal{F}}$-invariant.
+The definition of reduction of singularities gives that $D$ may intersect only non-dicritical components and that $\tilde{\mathcal{F}}$ is everywhere transverse to $D$.
+The $\pi$-image of a local leaf of $\mathcal{F}$ at each non-corner point of $D$ belongs to $\text{Sep}_{p}(\mathcal{F})$.
+Denote by $\text{Sep}_{p}(D)\subset\text{Sep}_{p}(\mathcal{F})$ the set of separatrices whose strict transforms by $\pi$ intersect the component $D\subset\mathcal{D}$.
+If $B\in\text{Sep}_{p}(D)$ with $D$ non-dicritical, $B$ is said to be isolated.
+Otherwise, it is said to be a dicritical separatrix.
+This engenders the decomposition $\text{Sep}_{p}(\mathcal{F})=\text{ISO}_{p}(\mathcal{F})\cup\text{Dic}_{p}(\mathcal{F})$, where notation is self evident.
+The set $\text{ISO}_{p}(\mathcal{F})$ is finite and contains all purely formal separatrices.
+It subdivides further in two classes: weak separatrices – those arising from the weak separatrices of saddle-nodes – and strong separatrices – corresponding to strong separatrices of saddle-nodes and separatrices of non-degenerate singularities.
+On the other hand, if $\text{Dic}_{p}(\mathcal{F})$ is a non-empty set then it is an infinite set of analytic separatrices of $\mathcal{F}$.
+A foliation $\mathcal{F}$ is said to be dicritical when $\text{Sep}_{p}(\mathcal{F})$ is infinite, which is equivalent to saying that $\text{Dic}_{p}(\mathcal{F})$ is non-empty.
+Otherwise, $\mathcal{F}$ is called non-dicritical.
+When we deal with a simple blow-up $\sigma$ at $p\in\mathbb{C}^{2}$, we say that $\sigma$ is dicritical with respect to $\mathcal{F}$, if the exceptional divisor $\sigma^{-1}(p)$ is not $\tilde{\mathcal{F}}$-invariant.
+Otherwise, $\sigma$ is called non-dicritical.
+Let $\mathcal{F}$ be a holomorphic foliation at $(\mathbb{C}^{2},p)$ and $B\in\text{Sep}_{p}(\mathcal{F})$ On a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we mark the reduction process of singularities of the pair $(\mathcal{F},B)$:
+(13) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
+where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ is the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(j+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$.
+Let
+(14) $\alpha_{j}:=\left\{\begin{array}[]{ll}\dim\frac{(\sigma_{j+2}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j+1)}}}{(\sigma_{j+1}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j)}}},&\hbox{ if $j=0,\ldots,N-2$}\\ \dim\frac{\Omega_{B^{(N)}}}{\sigma_{N}^{*}\Omega_{B^{(N-1)}}},&\hbox{ if $j=N-1$}\end{array}\right.$
+where for a curve $B$, $\Omega_{B}$ denotes the module of Kähler differentials on $B$.
+We note that $\alpha_{0}$ appears as a term in [19, Equation 4].
+Additionally, notice that $\alpha_{j}$ is the $\alpha_{0}$ for $B^{(j)}$.
+Moreover, we observe that $\alpha_{j}\geq 0$, and the equality holds when the curve $B^{(j)}$ is smooth.
+We set
+(15) $R_{p}(\mathcal{F},B):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
+where
+(16) $m_{q^{(j)}}(\mathcal{F}^{(j)})=\left\{\begin{array}[]{ll}\nu_{q^{(j)}}(\mathcal{F}^{(j)})+1&\hbox{\rm if $\sigma_{j+1}$ is dicritical}\\ \nu_{q^{(j)}}(\mathcal{F}^{(j)})&\hbox{\rm if $\sigma_{j+1}$ is non-dicritical}.\\ \end{array}\right.$
+Notice that for $j=0$ we get $q^{(0)}=p$, $\mathcal{F}^{(0)}=\mathcal{F}$ and the equality (16) for $j=0$ becomes:
+(17) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical},\\ \end{array}\right.$
+where $\sigma$ is a blow-up at the point $p$.
+The following lemma relates $R_{p}(\mathcal{F},B)$ to the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ (respectively $\mathcal{F}^{(N)}$) with respect to $B$ (respectively $B^{(N)}$).
+Lemma 3.1.
+With the above notations we get
+$R_{p}(\mathcal{F},B)=GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right].$
+Proof.
+From blow-up formula of the GSV index [4, p.30] we get
+(18) $GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})$ $=$ $GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})$
+$+$ $\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})),\,\,\,\,\hbox{\rm for }\,\,\,0\leq j\leq N-1.$
+Up substituting the equality (18) in (15), we get
+$R_{p}(\mathcal{F},B)$ $=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right]$
+$=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}))\right]$
+$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
+$=$ $\sum_{j=0}^{N-1}\left[GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})-GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})\right]$
+$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
+$=$ $GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)$
+$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right],$
+and the lemma follows. ∎
+For a germ of a holomorphic foliation $\mathcal{F}$ and a separatrix $B\in\text{Sep}_{p}(\mathcal{F})$ at $(\mathbb{C}^{2},p)$, we introduce the number $\text{DG}_{p}(\mathcal{F},B)$ as follows:
+(21) $\text{DG}_{p}(\mathcal{F},B)$ ::=:= $\left\{\begin{array}[]{ll}R_{p}(\mathcal{F},B)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})&\hbox{\rm if $B\in\text{ISO}_{p}(\mathcal{F})$}\\ R_{p}(\mathcal{F},B)&\hbox{\rm if $B\in\text{Dic}_{p}(\mathcal{F})$}.\\ \end{array}\right.$
+From (14), we deduce that $\alpha_{i}$ are analytic invariants, thus we have $\text{DG}_{p}(\mathcal{F},B)$ is also an analytic invariant depending of $\mathcal{F}$ and $B$.
+As a consequence of Lemma 3.1, we can establish $\text{DG}_{p}(\mathcal{F},B)$ in terms of the GSV-indices of the foliations $\mathcal{F}$ and $\mathcal{F}^{(N)}$ and some numbers depending of the resolution of the singularities of $B$.
 4. Proof of Theorem A
-Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
-(8) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
-Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (8):
-(9) $\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
-We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
-Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$).
-Denote by $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$.
-If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([15]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$.
-By [15, Claim 4.4] we have
-(10) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},$
-for $\nu_{p}(C)\geq 2$.
-We will need the following lemma.
+First of all, we will study the behaviour of $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ under a blow-up $\sigma$ at $p\in\mathbb{C}^{2}$.
+For this, we consider $\alpha_{0}$ as in (14) for $j=0$.
+We have the following lemma.
 Lemma 4.1.
-Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$.
-If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$, then
-(11) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
-Moreover, the equality holds if and only if $B$ is smooth.
+Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B\in\text{Sep}_{p}(\mathcal{F})$.
+If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by a blow-up $\sigma$ at $p$ then
+$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0},$
+where $q\in\tilde{B}\cap\sigma^{-1}(p)$, and $\alpha_{0}$ is as in (14).
 Proof.
-Suppose first $\nu_{p}(B)\geq 2$. By [5, Equality (1.2), p. 291] we get
+By [6, Equality (1.2), p. 291] we get
 (22) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
-where
-(13) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
-On the other hand by the behaviour under blow-up of the GSV index [3, p.30] we get
-(14) $GSV_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
-and by [7, Proposition 6.2] we have
+where $m_{p}(\mathcal{F})$ was introduced in (17).
+On the other hand by the behaviour under blow-up of the GSV index (see [4, p.30]) we get
+(23) $GSV_{q}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
+and after [8, Proposition 6.2] we have
 (24) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
-After [15, Equation (4)] we get:
-(16) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
-where $\alpha$ was defined in (10). From (15), (14), and (16) we deduce
-(17) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}+\alpha.$
+From [19, Equation (4)] we obtain
+(25) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha_{0},$
+where $\alpha_{0}$ was defined in (14). From (24), (23), and (25) we deduce
+(26) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha_{0}.$
 By (22), (23) and (26) it holds
-(18) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$ $=$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
--- $\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha.$
-Hence, if $B$ is a singular curve then the lemma follows by (10).
-Suppose now that $B$ is a non-singular curve.
-By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$.
-Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$.
-Hence $Q=xh$ for some formal power series $h(x,y)\in\mathbb{C}[[x,y]]$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$.
-Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
-(19) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,x)=\mu_{p}(\mathcal{F},B).$
-Moreover by (15) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
-Thus if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality.
-Moreover, if (11) is an equality, then $B$ is necessarily smooth.
-Indeed if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (18) we have that (11) is a strict inequality that is a contradiction.
-Hence the lemma follows.
-∎
+$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0}.$
+Remark 4.2.
+If $\mathcal{F}$ is a reduced foliation at $p\in\mathbb{C}^{2}$, and $B$ is a smooth separatrix through $p$, then, by Remark 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F}.B)=\tau_{p}(\mathcal{F},B)$.
+Therefore, according to [3, p.
+538], we have $\tau_{p}(\mathcal{F},B)=1$, if $p$ is a non-degenerate singularity, or if $p$ is a saddle-node singularity and $B$ is the strong separatrix; and $\tau_{p}(\mathcal{F},B)=k+1$, if $p$ is a saddle-node singularity and $B$ is the weak separatrix.
+Here $k\in\mathbb{Z}^{+}$ is a formal invariant in the case that $p$ is a saddle-node singularity of $\mathcal{F}$.
+In any cases, $\tau_{p}(\mathcal{F},B)>0$.
+We will now prove the main theorem of this paper.
+4.1. Proof of Theorem A
+Let $\mathcal{F}$ be a holomorphic foliation on $(\mathbb{C}^{2},p)$, and let $B$ be a separatrix of $\mathcal{F}$ at $p$.
+Then, on a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we marked the reduction process of the pair $(\mathcal{F},B)$:
+(27) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
+where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ coincide with the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$.
+Put $\mu^{(i)}=\mu_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$, for any $i\in\{0,\ldots,N\}$.
+Applying Lemma 4.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\overset{\sigma_{i+1}}{\longrightarrow}(\mathcal{F}^{(i)},B^{(i)})$, we get
+$3\mu^{(i)}-4\tau^{(i)}=3\mu^{(i+1)}-4\tau^{(i+1)}+\left[\nu_{q^{(i)}}(B^{(i)})(2\nu_{q^{(i)}}(B^{(i)})-m_{q^{(i)}}(\mathcal{F}^{(i)})-1)-4\alpha_{i}\right]$
+In particular
+$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ $=$ $3\mu^{(N)}-4\tau^{(N)}$
+$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
+$=$ $3\mu^{(N)}-4\tau^{(N)}+R_{p}(\mathcal{F},B),$
+where the last equality follows from (15).
+We have two possibilities: either $B\in\text{ISO}_{p}(\mathcal{F},B)$ or $B\in\text{Dic}_{p}(\mathcal{F})$.
+If $B\in\text{ISO}_{p}(\mathcal{F},B)$, there exists a component $D\subset\mathcal{D}$ with $D$ non-dicritical such that $\mathcal{F}^{(N)}$ has a reduced singularity at $q^{(N)}\in\mathcal{D}\cap B^{(N)}$.
+In this case, we have $\mu^{(N)}=\tau^{(N)}$ by Remark 4.2, and thus
+$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=-\tau^{(N)}+R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
+If $B\in\text{Dic}_{p}(\mathcal{F})$, then there exists a component $D\subset\mathcal{D}$ with $D$ dicritical such that $B^{(N)}$ is transversal to $D$ at $q^{(N)}$, and $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$.
+In this case, we get
+$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
+In both cases, the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ implies $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. $\Box$
+The following example illustrates that the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ in Theorem A cannot be removed.
+Example 4.3.
+Let $\mathcal{F}_{m}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
+$\omega_{m}=\left((2m+1)yx^{m+1}+my^{m+2}\right)dx+\left((1-m)xy^{m+1}-2mx^{m+2}\right)dy,$
+where $m\in\mathbb{Z}_{>0}$. Let $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$. We have that $B$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$ such that
+$GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}3&\text{if $m=1$}\\ -2m^{2}+4m+1&\text{ if $m>1$},\end{array}\right.$
+$\mu_{0}(B)=2m(2m-1)\,\,\,\,\,\,\,\,\,\,m\geq 1,$
+$\tau_{0}(B)=\left\{\begin{array}[]{cl}2&\text{if $m=1$}\\ 3m^{2}&\text{ if $m>1$},\end{array}\right.$
+$\mu_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ 2m^{2}+2m+1&\text{ if $m>1$}\end{array}\right.$
+and
+$\tau_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ m^{2}+4m+1&\text{if $m>1$}.\end{array}\right.$
+A straightforward calculation reveals that $\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}<\frac{4}{3}$ for $60<m<$ 6, and
+$\lim_{m\to+\infty}\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}=2.$
+On the other hand, we compute $\text{DG}_{0}(\mathcal{F}_{m},B)$ for $m\geq 1$. Note that $\nu_{0}(\mathcal{F}_{m})=m+2$, and $\nu_{0}(B)=2m$, and using [19, equation 4] we get
+$\alpha_{0}=\left\{\begin{array}[]{cl}1&\text{ if $m=1$}\\ m^{2}+m&\text{ if $m>1$}\end{array}\right.$
+After a blow-up at $0\in\mathbb{C}^{2}$, the strict transform of $B$ is a smooth curve, and $m_{0}(\mathcal{F}_{m})=m+2$.
+The reduction of singularities of the pair $(\mathcal{F}_{m},B)$ is achieved in $2m$ steps, and $B\in\text{Dic}_{0}(\mathcal{F}_{m})$.
+Thus
+$\text{DG}_{0}(\mathcal{F}_{1},B)$ $=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F})-1)-4\alpha_{0}\right]$
+$+\left[\nu_{q^{(1)}}(B^{(1)})(2\nu_{q^{(1)}}(B^{(1)})-m_{q^{(1)}}(\mathcal{F}^{(1)})-1)-4\alpha_{1}\right]$
+$=$ $\left[2(4-3-1)-4(1)\right]+\left[1(2-2-1)-4(0)\right]$
+$=$ $-5$
+and for $m>1$, we have
+$\text{DG}_{0}(\mathcal{F}_{m},B)$ $=$ $\sum_{j=0}^{2m-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}_{m})-1)-4\alpha_{j}\right]$
+$=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F}_{2m})-1)-4\alpha_{0}\right]$
+$+\left[\nu_{q^{(2m-1)}}(B^{(2m-1)})(2\nu_{q^{(2m-1)}}(B^{(2m-1)})-m_{q^{(2m-1)}}(\mathcal{F}^{(2m-1)}_{2m})-1)-4\alpha_{2m-1}\right]$
+$=$ $\left[2m\left(2(2m)-(m+2)-1\right)-4(m^{2}+m)\right]+\left[1(2-2-1)-4(0)\right]$
+$=$ $1\displaystyle 2m^{2}-10m-$ 1 .
+Therefore, $\text{DG}_{0}(\mathcal{F}_{m},B)\geq 0$ if and only if $m\geq 6$. This shows that the hypothesis of Theorem A is optimal.
+5. Applications
+In this section, from the proof of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
 Corollary 5.1.
-Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$.
-If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) at the point $q$ after a composition of $\ell$ blows-up then
-(20) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
-Proof.
-It follows applying Lemma 3.1 $\ell$ times. ∎
-Remark 3.3.
-By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
-(21) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
-In addition, by (19) if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
-Remark 3.4.
-If $\mathcal{F}$ is a non-dicritical second type foliation at $p\in\mathbb{C}^{2}$, and $B$ is the only separatrix of $\mathcal{F}$, then $\mathcal{F}$ is a generalized curve foliation.
-In fact, assume by contradiction that $\mathcal{F}$ admits a saddle-node singularity in its reduction process with strong separatrix on the exceptional divisor.
-Without lost of generality, we can suppose that $\mathcal{F}$ is a reduced foliation after a blow-up $\sigma$ at $p$.
-Since $\mathcal{F}$ is of second type, after a blow-up, the strong separatrix of the strict transform $\tilde{\mathcal{F}}$ of $\mathcal{F}$ through $q\in\sigma^{-1}(p)$ is contained in $E=\sigma^{-1}(p)$, thus the Camacho-Sad index is $CS_{q}(\tilde{\mathcal{F}},E)=0$, but it is a contradiction with Camacho-Sad’s formula
-$CS_{q}(\tilde{\mathcal{F}},E)=E\cdot E=-1.$
-Proof of Theorem A: If $B$ is a non-singular curve then, by Remark 3.3, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$ and the theorem follows.
-Let us suppose now that $B$ is a singular separatrix, that is $\nu_{p}(B)\geq 2$.
-After Corollary 3.2 we can suppose, without lost of generality that the number of blows-up $n$ needed in the process of the reduction of the singularity of $B$ is one.
-In particular, after a blow-up $\sigma$ the strict transform $\tilde{B}$ of $B$ is smooth, that is the algebraic multiplicity of $\tilde{B}$ is one.
-Applying Lemma 3.1 and after (14) we get
-(22) $[3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})]-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
-Since $\tilde{B}$ is smooth, from (22) and the equality $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{q}(\tilde{\mathcal{F}},\tilde{B})$ (see Remark 3.3) we obtain
-(23) $-\tau_{q}(\tilde{\mathcal{F}},\tilde{B})-[3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)]>\nu_{p}(B)(m_{p}(\mathcal{F})-\nu_{p}(B)).$
-Moreover, by the definition of $m_{p}(\mathcal{F})$ and after [11, Proposition 3.3], if $\mathcal{B}=B+\mathcal{B}^{\prime}$ is a balanced divisor of separatrices of $\mathcal{F}$ (for some divisor $\mathcal{B}^{\prime}$, if $B$ is the only separatrix of $\mathcal{F}$ we consider for $\mathcal{B}^{\prime}$ the empty divisor) then we have
-$m_{p}(\mathcal{F})-\nu_{p}(B)=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{B}^{\prime})-1+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical}\\ \nu_{p}(\mathcal{B}^{\prime})+\xi_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is dicritical},\end{array}\right.$
-where $\xi_{p}(\mathcal{F})$ is the tangency excess of $\mathcal{F}$ (see for example [7, Definition 2.3]).
-So $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$ except in the case that $\mathcal{F}$ is a foliation of the second type ($\xi_{p}(\mathcal{F})=0$) and $B$ is its only separatrix ($\nu_{p}(\mathcal{B}^{\prime})=0$) in which case we get $m_{p}(\mathcal{F})-\nu_{p}(B)=-1$.
-In the case $m_{p}(\mathcal{F})-\nu_{p}(B)\geq 0$, from (23) we obtain $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows.
-Otherwise, assume that $\mathcal{F}$ is of the second type and that $B$ is its only separatrix.
-By Remark 3.4, $\mathcal{F}$ is a generalized curve foliation, so $GSV_{p}(\mathcal{F},B)=0$.
-Moreover, since $\tilde{B}$ is smooth then the left-term of (11) is zero, so the theorem follows by Lemma (3.1) (since $B$ is singular the inequality (11) is strict) and Remark 3.3 ($\tilde{B}$ is smooth).
-$\Box$
-From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [15]:
-Corollary 3.5.
-Let $f(x,y)\in\mathbb{C}[[x,y]]$ an irreducible power series. Put $C:f(x,y)=0$ the branch defined by $f$. Then
+Let $f(x,y)\in\mathbb{C}\{x,y\}$ an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
 $\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
 Proof.
-It is a consequence of Remark 2.1 and of applying Theorem A to foliation $\mathcal{F}:df=0$. ∎
-Remark 3.6.
-In [15] Wang proved that the number $3\mu_{p}(B)-4\tau_{p}(B)$ is a non-negative number, non-decreasing under blowups and strictly increasing unless the curve $B$ is non-singular.
-The result of Wang is an special case of Lemma 3.1, taking $\mathcal{F}=df$ where $B$ is given by $f(x,y)=0$.
-Lemma 3.1 and Theorem A generalizes this for the number $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$.
-Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
+Consider the singular holomorphic foliation $\mathcal{F}:\omega=df$. In this situation, $C\in\text{ISO}_{0}(\mathcal{F})$. To continue, we compute
+$\text{DG}_{0}(\mathcal{F},C)=R_{0}(\mathcal{F},C)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)}),$
+where
+$R_{0}(\mathcal{F},C):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(2\nu_{q^{(j)}}(C^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
+and $C^{(j+1)}$ denotes the strict transform of $C^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $C^{(N)}$ is the normalization of $C$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in C^{(N)}$ as in the proof of Theorem A.
+Since $C^{(N)}$ is smooth, we have $\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})=GSV_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})$ by Remark 4.2, and from Lemma 3.1, we deduce
+$\text{DG}_{0}(\mathcal{F},C)=-GSV_{0}(\mathcal{F},C)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right]$
+Now, since $\mathcal{F}$ is a curve generalized non-dicritical foliation at $0\in\mathbb{C}^{2}$, and $\text{Sep}_{0}(\mathcal{F})=C$, we have $GSV_{0}(\mathcal{F},C)=0$ by [3, p.
+528].
+Moreover $\mu_{0}(\mathcal{F},C)=\mu(C)$ and $\tau_{0}(\mathcal{F},C)=\tau(C)$ by Remark 2.1.
+Hence
+$\text{DG}_{0}(\mathcal{F},C)=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right].$
+Finally, applying successively [19, Claim 4.4] in each step of the resolution of singularities of $C$, and given that the corresponding value $\mathcal{D}_{min}$ of [19, Claim 4.4] in the $j$-th step of the resolution of $C$ is less or equal to $\alpha_{j}$, we get $\text{DG}_{0}(\mathcal{F},C)<0$ implying
+$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}$
+by Theorem A. ∎
+Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
 Corollary 5.2.
-Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ be a balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
-(24) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
+Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$.
+Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$.
+If $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$ with $\text{DG}_{p}(\mathcal{F},B_{j})<0$ for all $1\leq j\leq\ell$, then
+(28) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}<\frac{4}{3}.$
 Proof.
-By (6) we have $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}$.
-Hence we will prove that
-(25) $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}<\frac{4}{3}.$
-Suppose that $\mathcal{B}_{0}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
+Suppose that $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
 $\sum_{j=1}^{\ell}\left(3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})\right)\leq-\ell<-\ell+1.$
@@
 so
-$\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
-Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B}_{0})}{T_{p}(\mathcal{F},\mathcal{B}_{0})-\ell+1}<\frac{4}{3}.$
+$\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
+Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
 The following example illustrates Corollary 5.2.
@@
 Let $\mathcal{F}:\omega=(2x^{7}+5y^{5})dx-xy^{2}(5y^{2}+3x^{5})dy$, whose leaves are contained in the connected components of the curves $C_{\lambda,\zeta}:\zeta(y^{5}-x^{7}+x^{5}y^{3})-\lambda x^{5}=0$, where $(\zeta,\lambda)\neq(0,0)$.
-Let $C=\{x=0\}$, $\mathcal{B}_{0}=C+C_{0,1}+C_{1,1}$, $\mathcal{B}_{\infty}=C_{2,1}$, then $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices for $\mathcal{F}$.
-We have
-$\mu_{0}(\mathcal{F},\mathcal{B}_{0})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{0,1})+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+29+21-2+1=54,$
-$\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})=\mu_{0}(\mathcal{F},C_{2,1})=21,$
-and
-$T_{0}(\mathcal{F},\mathcal{B}_{0})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{0,1})+\tau_{0}(\mathcal{F},C_{1,1})=5+28+21=54.$
-Hence
-$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},\mathcal{B}_{0})-\mu_{0}(\mathcal{F},\mathcal{B}_{\infty})+1=54-21+1=34,$
-and
-$\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B}_{0})}=\frac{34}{54}<\frac{4}{3}.$
-We conclude the paper with the following question:
-Question: Is the inequality (24) true if instead of $T_{p}(\mathcal{F},\mathcal{B}_{0})$ we write $\tau_{p}(\mathcal{F},\mathcal{B}_{0})$?
+Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$.
+We have $GSV_{0}(\mathcal{F},C)=5$, $GSV_{0}(\mathcal{F},C_{1,1})=5$, $\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$, hence $\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B})}=\frac{25}{26}<\frac{4}{3}$.
+We conclude the paper with a question:
+Question 1: Is the inequality (28) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?

### v3

@@
 Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$.
-Under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we prove the Dimca-Greuel type inequality $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<4/3$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$.
-As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes but it is in line with the idea developed by Wang.
+We prove the following Dimca-Greuel type inequality $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$, $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$, and $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$.
+As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes, but it is in line with the idea developed by Wang.
 Key words and phrases:
@@
 Primary 32S65 - 32M25
-Acknowledgments.
-The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done (Spanish grant PID2019-105896GB-I00 funded).
-This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979.
 1. Introduction
@@
 Is clear from their definitions that $\tau(C)\leq\mu(C)$.
-In [7, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity.
+In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity.
 The question was affirmatively answered in [1], [10], and [17] for the irreducible case and in [2] for the reduced case.
 It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs.
-More precisely, if $\mathcal{F}$ is a germ of singular foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}[[x,y]]/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$.
-The algebraic multiplicity $\nu_{p}(\mathcal{F})$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
-Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [10], [12]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension $\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}[[x,y]]/(P,Q,f)$.
+More precisely, if $\mathcal{F}$ is a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}\{x,y\}/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$.
+Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9, p.
+4], [11, Definition 3.1]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension
+$\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
 If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
@@
 equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$.
-If the foliation $\mathcal{F}$ is of the second type (see [10, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [9, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [15] where the general case of hypersurfaces is studied).
+If the foliation $\mathcal{F}$ is of the second type (see [9, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied).
 It is natural to wonder if we can improve the bound $2$ for the ratio (1).
-In general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [8, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
+We note that, in general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
 $\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
 then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
-However, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and we put in the denominator $\tau_{p}(\mathcal{F},B)$, then, under certain conditions on the reduction of singularities of $\mathcal{F}$ and $B$, we will obtain a bound of the Dimca-Greuel type for foliations.
-More specifically, in Foliation theory, two type of separatrices of $\mathcal{F}$ are distinguished: isolated separatrices $\text{ISO}_{p}(\mathcal{F})$ and dicritical separatrices $\text{Dic}_{p}(\mathcal{F})$.
-Isolated separatrices of $\mathcal{F}$ correspond to separatrices that are images, by the reduction process of singularities $\pi$ of $\mathcal{F}$, of separatrices of non-degenerate singularities, or weak or strong separatrices of saddle-nodes singularities in a reduction of singularities of $\mathcal{F}$.
-On the other hand, dicritical separatrices correspond to separatrices that are images, by $\pi$, of separatrices that intersect transversally with a dicritical divisor in the last step of the reduction of singularities of $\mathcal{F}$.
-For more details about these definitions, see Section 3.
-In order to establish an inequality of Dimca-Greuel type, it is necessary to introduce a number $\text{DG}_{p}(\mathcal{F},B)$ that depends on the reduction of singularities of the pair $(\mathcal{F},B)$ and on the type of separatrix that $B$ is, see Definition 21.
-We will see that the negativity of this number is crucial in order to obtain an inequality of the Dimca-Greuel type for foliations.
-Our main theorem is
+On the other hand, following Camacho-Lins Neto-Sad [4, Section 4], we can consider the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition).
+We note that if $B=\{f(x,y)=0\}$, and $\mathcal{F}:df=0$, then $\mu_{p}(\mathcal{F},B)$ also coincides with the Milnor number $\mu(B)$ of $B$ (see Remark 2.1).
+Therefore, we could think to studying if the inequality
+$03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0
+holds. But, for the germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},0)$ defined by
+(2) $\omega=y\left(13x^{7}+6y^{7}\right)dx-x\left(5y^{7}+12x^{7}\right)dy,$
+and the separatrix
+(3) $B:x^{13}+x^{6}y^{7}+y^{12}=0,$
+we have $\mu_{0}(\mathcal{F},B)=85$ and $\tau_{0}(\mathcal{F},B)=61$. Thus
+$113\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61=255-244=$ 11 .
+In order to establish an inequality of the Dimca-Greuel’s type for foliations, it is necessary to add the Gómez-Mont–Seade–Verjovsky index (in short GSV-index) (see [12]) as stating our main result.
 Theorem A.
-Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. If $\text{DG}_{p}(\mathcal{F},B)<0$, then
-(2) $\frac{\mu_{p}(\mathcal{F},B)}{\tau_{p}(\mathcal{F},B)}<\frac{4}{3}.$
-We will prove Theorem A in Section 4. In Example 4.3, we will show that we cannot removed the hypothesis over $\text{DG}_{p}(\mathcal{F},B)$.
-As a consequence of Theorem A we deduce (see Corollary 5.1) the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
+Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
+(4) $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$
+where $GSV_{p}(\mathcal{F},B)$ denotes the GSV-index of $\mathcal{F}$ with respect to $B$. Moreover, the equality holds if and only if $B$ is smooth.
+We will prove Theorem A in Section 3. For the foliation given in (2), and $B$ as in (3), we have $GSV_{0}(\mathcal{F},B)=-47$. Hence,
+$03\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)+GSV_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61+(-47)=-36<$ 0
+verifying Theorem A.
+As a consequence of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17] (see Corollary 5.1).
 2. Multiplicity of a foliation along a divisor of separatrices
-Throughout this note $\mathcal{F}$ denotes a germ of a singular (holomorphic or formal) foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a (holomorphic or formal) 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}[[x,y]]$ are coprime; or by its dual vector field
-(3) $v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
-Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}[[x,y]]$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
-We denote by $\text{Sep}_{p}(\mathcal{F})$ the set of all separatrices of $\mathcal{F}$ at $p$.
-Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by a vector field $v$ as in (3) and $B\in\text{Sep}_{p}(\mathcal{F})$.
-Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, we can consider the multiplicity of $\mathcal{F}$ along $B$ at $p$ defined by
+Throughout this note $\mathcal{F}$ denotes a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$, given in local coordinates $(x,y)$ centered at $p$ by a 1-form $\omega:=P(x,y)dx+Q(x,y)dy$, where $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$ are coprime; or by its dual vector field
+$v:=-Q(x,y)\frac{\partial}{\partial x}+P(x,y)\frac{\partial}{\partial y}.$
+The algebraic multiplicity $\nu_{p}(\mathcal{F})$ of $\mathcal{F}$ is the minimum of the orders $\text{ord}_{p}(P)$ and $\text{ord}_{p}(Q)$.
+Remember that a plane curve germ $f(x,y)=0$ is an $\mathcal{F}-$invariant curve if $\omega\wedge df=(f.h)dx\wedge dy$, for some $h(x,y)\in\mathbb{C}\{x,y\}$ and a separatrix of $\mathcal{F}$ is an irreducible $\mathcal{F}$-invariant curve.
+Let $\mathcal{F}$ be a germ of a singular foliation at $(\mathbb{C}^{2},p)$ induced by the vector field $v$ and $B$ be a separatrix of $\mathcal{F}$.
+Let $\gamma:(\mathbb{C},0)\to(\mathbb{C}^{2},p)$ be a primitive parametrization of $B$, the multiplicity of $\mathcal{F}$ along $B$ at $p$ is by definition
 (5) $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}\theta(t),$
-where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [5, p.
+where $\theta(t)$ is the unique vector field at $(\mathbb{C},0)$ such that $\gamma_{*}\theta(t)=v\circ\gamma(t)$, see for instance [4, p.
 159].
@@
 By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$.
-In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [14, Equation (2.2), p.
+In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p.
 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
 Hence, if we write $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ where $\mathcal{B}_{0}$ and $\mathcal{B}_{\infty}$ are effective divisors we get
-(7) $\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
+$\mu_{p}(\mathcal{F},\mathcal{B})=\mu_{p}(\mathcal{F},\mathcal{B}_{0})-\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})+1.$
 Since $\mu_{p}(\mathcal{F},\mathcal{B}_{\infty})\geq 1$ we have
-(8) $\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
+$\mu_{p}(\mathcal{F},\mathcal{B}_{0})\geq\mu_{p}(\mathcal{F},\mathcal{B}).$
 Denote by $GSV_{p}(\mathcal{F},\mathcal{B})$ the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ (GSV-index) with respect to the effective divisor $\mathcal{B}$ of separatrices of $\mathcal{F}$.
@@
 Let $B:f(x,y)=0$ be a separatrix of $\mathcal{F}$ and let $\mathcal{G}_{f}$ be the hamiltonian foliation defined by $df=0$. From (5) and (6), we have
-(9) $\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
+$\mu_{p}(\mathcal{G}_{f},B)=\begin{cases}\text{ord}_{t}\partial_{y}{f}(\gamma(t))-\text{ord}_{t}x(t)+1&\text{if $x(t)\neq 0$}\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \text{ord}_{t}\partial_{x}{f}(\gamma(t))-\text{ord}_{t}y(t)+1&\text{if $y(t)\neq 0$},\end{cases}$
 where $\gamma(t)=(x(t),y(t))$ is a Puiseux parametrization of $B$.
-By [8, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$.
-But, since $B$ is irreducible, after [8, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [8, Proposition 6.2], we obtain
+By [7, Proposition 5.7], we get $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(\mathcal{G}_{f},B)$.
+But, since $B$ is irreducible, after [7, Proposition 4.7], $\mu_{p}(\mathcal{G}_{f},B)=\mu_{p}(B)$ and so that $\mu_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)+\mu_{p}(B).$ Hence, according to [7, Proposition 6.2], we obtain
 (8) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
-Consequently when $B$ is quasihomogeneous by ([17], Satz p.123 a) and d)] and (10) or by [20, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
-3. Reduction of singularities of foliations and their separatrices
-We say that $p\in\mathbb{C}^{2}$ is a reduced singularity for $\mathcal{F}$ if the linear part $\text{D}v(p)$ of the vector field $v$ in (3) is non-zero and has eigenvalues $\lambda_{1},\lambda_{2}\in\mathbb{C}$ fitting in one of the cases:
-  1. (i)
-    $\lambda_{1}\lambda_{2}\neq 0$ and $\lambda_{1}\lambda_{2}\not\in\mathbb{Q}^{+}$ (non-degenerate);
-  2. (ii)
-    $\lambda_{1}\neq 0$ and $\lambda_{2}\neq 0$ (saddle-node singularity).
-In the case $(i)$, there is a system of coordinates $(x,y)$ in which $\mathcal{F}$ is defined by the equation
-(11) $\omega=x(\lambda_{1}+a(x,y))dy-y(\lambda_{2}+b(x,y))dx,$
-where $a(x,y),b(x,y)\in{\mathbb{C}}[[x,y]]$ are non-units, so that $\text{Sep}_{p}(\mathcal{F})$ is formed by two transversal analytic branches given by $\{x=0\}$ and $\{y=0\}$.
-In the case $(ii)$, up to a formal change of coordinates, the saddle-node singularity is given by a 1-form of the type
-(12) $\omega=x^{k+1}dy-y(1+\lambda x^{k})dx,$
-where $\lambda\in\mathbb{C}$ and $k\in\mathbb{Z}^{+}$ are invariants after formal changes of coordinates (see [16, Proposition 4.3]).
-The curve $\{x=0\}$ is an analytic separatrix, called strong separatrix, whereas $\{y=0\}$ corresponds to a possibly formal separatrix, called weak separatrix.
-The integer $k+1>1$ is called tangency index of $\mathcal{F}$ with respect to the weak separatrix.
-Let $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ be a composition of blow-ups maps.
-The divisor $\mathcal{D}$ is a finite union of components which are embedding projective lines, crossing normally at corners.
-If $\mathcal{F}$ is the foliation given by the 1-form $\omega$, we denote ~$\tilde{\mathcal{F}}=\pi^{*}\mathcal{F}$ the strict transform of $\mathcal{F}$, the germ of foliation on $(\tilde{X},\mathcal{D})$ defined locally by $\pi^{*}(\omega)$, after cancelling one-dimensional singular components.
-For a uniform analysis, we include the possibility of $\pi$ being the identity map and, abusing notation, we set in this case $\tilde{X}=\mathbb{C}^{2}$, $\mathcal{D}=\{p\}$ and $\tilde{\mathcal{F}}=\mathcal{F}$.
-With respect to the divisor $\mathcal{D}$, the foliation $\tilde{\mathcal{F}}$ at a point $q\in\mathcal{D}$ can be:
-  • Regular, if there are local analytic coordinates $(x,y)$ at $q$ such that $\mathcal{D}\subset\{xy=0\}$ and ~$\tilde{\mathcal{F}}:dx=0$.
-  • Singular; if it is not regular.
-  • Reduced, if $q$ is a reduced singularity for $\tilde{\mathcal{F}}$ and $\mathcal{D}\subset\text{Sep}_{q}(\tilde{\mathcal{F}})$.
-When $\mathcal{D}=\{p\}$, these notions coincide with the ordinary concepts of regular point, singular point and reduced singularity.
-We say that $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ is a reduction of singularities for $\mathcal{F}$ if all points $q\in\mathcal{D}$ are either regular or reduced singularities.
-There always exists a reduction of singularities [18].
-Besides, there exists a minimal one, in the sense that it factorizes any other reduction of singularities by an additional sequence of blow-ups.
-All along this note, reductions of singularities are supposed to be minimal.
-Given a germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},p)$ with a fixed reduction process $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},0)$ for $\mathcal{F}$, a component $D\subset\mathcal{D}$ can be:
-  • Non-dicritical, if $D$ is $\tilde{\mathcal{F}}$-invariant.
-In this case $D$ contains a finite number of reduced singularities.
-Each non-corner singularity carries a separatrix transversal to $D$, whose projection by $\pi$ is a curve in $\text{Sep}_{p}(\mathcal{F})$.
-  • Dicritical, if $D$ is not $\tilde{\mathcal{F}}$-invariant.
-The definition of reduction of singularities gives that $D$ may intersect only non-dicritical components and that $\tilde{\mathcal{F}}$ is everywhere transverse to $D$.
-The $\pi$-image of a local leaf of $\mathcal{F}$ at each non-corner point of $D$ belongs to $\text{Sep}_{p}(\mathcal{F})$.
-Denote by $\text{Sep}_{p}(D)\subset\text{Sep}_{p}(\mathcal{F})$ the set of separatrices whose strict transforms by $\pi$ intersect the component $D\subset\mathcal{D}$.
-If $B\in\text{Sep}_{p}(D)$ with $D$ non-dicritical, $B$ is said to be isolated.
-Otherwise, it is said to be a dicritical separatrix.
-This engenders the decomposition $\text{Sep}_{p}(\mathcal{F})=\text{ISO}_{p}(\mathcal{F})\cup\text{Dic}_{p}(\mathcal{F})$, where notation is self evident.
-The set $\text{ISO}_{p}(\mathcal{F})$ is finite and contains all purely formal separatrices.
-It subdivides further in two classes: weak separatrices – those arising from the weak separatrices of saddle-nodes – and strong separatrices – corresponding to strong separatrices of saddle-nodes and separatrices of non-degenerate singularities.
-On the other hand, if $\text{Dic}_{p}(\mathcal{F})$ is a non-empty set then it is an infinite set of analytic separatrices of $\mathcal{F}$.
-A foliation $\mathcal{F}$ is said to be dicritical when $\text{Sep}_{p}(\mathcal{F})$ is infinite, which is equivalent to saying that $\text{Dic}_{p}(\mathcal{F})$ is non-empty.
-Otherwise, $\mathcal{F}$ is called non-dicritical.
-When we deal with a simple blow-up $\sigma$ at $p\in\mathbb{C}^{2}$, we say that $\sigma$ is dicritical with respect to $\mathcal{F}$, if the exceptional divisor $\sigma^{-1}(p)$ is not $\tilde{\mathcal{F}}$-invariant.
-Otherwise, $\sigma$ is called non-dicritical.
-Let $\mathcal{F}$ be a holomorphic foliation at $(\mathbb{C}^{2},p)$ and $B\in\text{Sep}_{p}(\mathcal{F})$ On a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we mark the reduction process of singularities of the pair $(\mathcal{F},B)$:
-(13) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
-where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ is the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(j+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$.
-Let
-(14) $\alpha_{j}:=\left\{\begin{array}[]{ll}\dim\frac{(\sigma_{j+2}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j+1)}}}{(\sigma_{j+1}\circ\ldots\circ\sigma_{N})^{*}\Omega_{B^{(j)}}},&\hbox{ if $j=0,\ldots,N-2$}\\ \dim\frac{\Omega_{B^{(N)}}}{\sigma_{N}^{*}\Omega_{B^{(N-1)}}},&\hbox{ if $j=N-1$}\end{array}\right.$
-where for a curve $B$, $\Omega_{B}$ denotes the module of Kähler differentials on $B$.
-We note that $\alpha_{0}$ appears as a term in [19, Equation 4].
-Additionally, notice that $\alpha_{j}$ is the $\alpha_{0}$ for $B^{(j)}$.
-Moreover, we observe that $\alpha_{j}\geq 0$, and the equality holds when the curve $B^{(j)}$ is smooth.
-We set
-(15) $R_{p}(\mathcal{F},B):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
+Consequently when $B$ is quasihomogeneous by ([15], Satz p.123 a) and d)] and (8) or by [18, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
+3. Proof of Theorem A
+Let $C$ be an irreducible plane curve of multiplicity $\nu_{p}(C)=n$. Denote by $\tilde{C}$ the strict transform of $C$ by a blow-up and $\bar{C}$ the normalization of $C$. We have natural morphisms
+(9) $\pi:\bar{C}\buildrel\rho\over{\longrightarrow}\tilde{C}\buildrel\sigma\over{\longrightarrow}C.$
+Consider $(C,p)\buildrel i\over{\longrightarrow}(\mathcal{C},x_{0})\buildrel\phi\over{\longrightarrow}(T,t_{0})$ the miniversal deformation of $C$. We extend to $\mathcal{C}$ the morphisms in (9):
+$\pi:\bar{\mathcal{C}}\buildrel\rho\over{\longrightarrow}\tilde{\mathcal{C}}\buildrel\sigma\over{\longrightarrow}\mathcal{C}.$
+We have two other natural morphisms $\tilde{C}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{C}\buildrel\bar{\phi}\over{\longrightarrow}T$ and its respectively extensions to the deformation, that is $\tilde{\mathcal{C}}\buildrel\tilde{\phi}\over{\longrightarrow}T$ and $\bar{\mathcal{C}}\buildrel\bar{\phi}\over{\longrightarrow}T$.
+Let $\Omega_{\mathcal{C}/T}$ (respectively, $\Omega_{\tilde{\mathcal{C}}/T}$, respectively, $\Omega_{\bar{\mathcal{C}}/T}$ ) denote the $\mathcal{O}_{\mathcal{C}}$-module of relative Kähler differentials of $\mathcal{C}$ over $T$ (respectively, the $\mathcal{O}_{\tilde{\mathcal{C}}}$-module of relative Kähler differentials of $\tilde{\mathcal{C}}$ over $T$, respectively the $\mathcal{O}_{\bar{\mathcal{C}}}$-module of relative Kähler differentials of $\bar{\mathcal{C}}$ over $T$).
+Put $G:=\rho^{*}\Omega_{\tilde{\mathcal{C}}/T}/\pi^{*}\Omega_{{\mathcal{C}}/T}$.
+If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([17]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$.
+By [17, Claim 4.4] we have
+(10) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},\,\,\,\,\text{for}\,\,\,\,\nu_{p}(C)\geq 2.$
+We will need the following lemma.
+Lemma 3.1.
+Let $\mathcal{F}$ be a germ of a singular foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$.
+If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by $\sigma$ at the point $q$, then
+(11) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})\geq 3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B).$
+Moreover, the equality holds if and only if $B$ is smooth.
+Proof.
+Suppose first $\nu_{p}(B)\geq 2$. By [5, Equality (1.2), p. 291] we have
+(12) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
 where
-(16) $m_{q^{(j)}}(\mathcal{F}^{(j)})=\left\{\begin{array}[]{ll}\nu_{q^{(j)}}(\mathcal{F}^{(j)})+1&\hbox{\rm if $\sigma_{j+1}$ is dicritical}\\ \nu_{q^{(j)}}(\mathcal{F}^{(j)})&\hbox{\rm if $\sigma_{j+1}$ is non-dicritical}.\\ \end{array}\right.$
-Notice that for $j=0$ we get $q^{(0)}=p$, $\mathcal{F}^{(0)}=\mathcal{F}$ and the equality (16) for $j=0$ becomes:
-(17) $m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is non-dicritical},\\ \end{array}\right.$
-where $\sigma$ is a blow-up at the point $p$.
-The following lemma relates $R_{p}(\mathcal{F},B)$ to the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ (respectively $\mathcal{F}^{(N)}$) with respect to $B$ (respectively $B^{(N)}$).
-Lemma 3.1.
-With the above notations we get
-$R_{p}(\mathcal{F},B)=GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right].$
-Proof.
-From blow-up formula of the GSV index [4, p.30] we get
-(18) $GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})$ $=$ $GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})$
-$+$ $\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})),\,\,\,\,\hbox{\rm for }\,\,\,0\leq j\leq N-1.$
-Up substituting the equality (18) in (15), we get
-$R_{p}(\mathcal{F},B)$ $=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right]$
-$=$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}))\right]$
-$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
-$=$ $\sum_{j=0}^{N-1}\left[GSV_{q^{(j+1)}}(\mathcal{F}^{(j+1)},B^{(j+1)})-GSV_{q^{(j)}}(\mathcal{F}^{(j)},B^{(j)})\right]$
-$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right]$
-$=$ $GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})-GSV_{p}(\mathcal{F},B)$
-$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(\nu_{q^{(j)}}(B^{(j)})-1)-4\alpha_{j}\right],$
-and the lemma follows. ∎
-For a germ of a holomorphic foliation $\mathcal{F}$ and a separatrix $B\in\text{Sep}_{p}(\mathcal{F})$ at $(\mathbb{C}^{2},p)$, we introduce the number $\text{DG}_{p}(\mathcal{F},B)$ as follows:
-(21) $\text{DG}_{p}(\mathcal{F},B)$ ::=:= $\left\{\begin{array}[]{ll}R_{p}(\mathcal{F},B)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})&\hbox{\rm if $B\in\text{ISO}_{p}(\mathcal{F})$}\\ R_{p}(\mathcal{F},B)&\hbox{\rm if $B\in\text{Dic}_{p}(\mathcal{F})$}.\\ \end{array}\right.$
-From (14), we deduce that $\alpha_{i}$ are analytic invariants, thus we have $\text{DG}_{p}(\mathcal{F},B)$ is also an analytic invariant depending of $\mathcal{F}$ and $B$.
-As a consequence of Lemma 3.1, we can establish $\text{DG}_{p}(\mathcal{F},B)$ in terms of the GSV-indices of the foliations $\mathcal{F}$ and $\mathcal{F}^{(N)}$ and some numbers depending of the resolution of the singularities of $B$.
-4. Proof of Theorem A
-First of all, we will study the behaviour of $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ under a blow-up $\sigma$ at $p\in\mathbb{C}^{2}$.
-For this, we consider $\alpha_{0}$ as in (14) for $j=0$.
-We have the following lemma.
-Lemma 4.1.
-Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B\in\text{Sep}_{p}(\mathcal{F})$.
-If $\tilde{\mathcal{F}}$ (respectively $\tilde{B}$) is the strict transform of $\mathcal{F}$ (respectively of $B$) by a blow-up $\sigma$ at $p$ then
-$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0},$
-where $q\in\tilde{B}\cap\sigma^{-1}(p)$, and $\alpha_{0}$ is as in (14).
-Proof.
-By [6, Equality (1.2), p. 291] we get
-(22) $\mu_{q}(\tilde{\mathcal{F}},\tilde{B})=\mu_{p}(\mathcal{F},B)-\nu_{p}(B)(m_{p}(\mathcal{F})-1),$
-where $m_{p}(\mathcal{F})$ was introduced in (17).
-On the other hand by the behaviour under blow-up of the GSV index (see [4, p.30]) we get
+$m_{p}(\mathcal{F})=\left\{\begin{array}[]{ll}\nu_{p}(\mathcal{F})+1&\hbox{\rm if$\sigma$ is dicritical}\\ \nu_{p}(\mathcal{F})&\hbox{\rm if $\sigma$ is nondicritical}.\\ \end{array}\right.$
+On the other hand by the behavior under blow-up of the GSV index [3, p. 30] we get
 (13) $GSV_{q}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\mathcal{F},B)+\nu_{p}(B)(\nu_{p}(B)-m_{p}(\mathcal{F})),$
-and after [8, Proposition 6.2] we have
+and after [7, Proposition 6.2] we have
 (14) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
-From [19, Equation (4)] we obtain
-(25) $\tau_{p}(B)-\tau_{q}(\tilde{B})=\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha_{0},$
-where $\alpha_{0}$ was defined in (14). From (24), (23), and (25) we deduce
-(26) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha_{0}.$
-By (22), (23) and (26) it holds
-$3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})=3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)-\nu_{p}(B)(2\nu_{p}(B)-m_{p}(\mathcal{F})-1)+4\alpha_{0}.$
+From [17, Equation (4)] we obtain
+(15) $\tau_{p}(B)-\tau_{q}(\tilde{B})\geq\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
+where $\alpha$ was defined in (10). From (14), (13), and (15) we deduce
+(16) $\tau_{q}(\tilde{\mathcal{F}},\tilde{B})\leq\tau_{p}(\mathcal{F},B)+\dfrac{\nu_{p}(B)[\nu_{p}(B)+1-2m_{p}(\mathcal{F})]}{2}-\alpha.$
+After (12), (13) and (16) we get
+(17) $3\mu_{q}(\tilde{\mathcal{F}},\tilde{B})-4\tau_{q}(\tilde{\mathcal{F}},\tilde{B})+GSV_{q}(\tilde{\mathcal{F}},\tilde{B})$$\geq$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)$
+$+$ $4\alpha-\nu_{p}(B)(\nu_{p}(B)-1)$
+$>$ $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B),$
+where the last inequality follows by (10), since $\nu_{p}(B)\geq 2.$
+Suppose now that $B$ is a non-singular curve.
+By a change of coordinates, if necessary, we can suppose that $B$ is given by $x=0$.
+Since $B$ is a separatrix of $\mathcal{F}:\omega=Pdx+Qdy$ then $\omega\wedge dx=Qdx\wedge dy$.
+Hence $Q=xh$ for some convergent power series $h(x,y)\in\mathbb{C}\{x,y\}$ such that $h(0,y)\not=0$ and $\mathcal{F}:\omega=P(x,y)dx+xh(x,y)dy$.
+Since $\gamma(t)=(0,t)$ is a parametrization of $B$, then $\mu_{p}(\mathcal{F},B)=\text{ord}_{t}P(0,t)$ and
+(18) $\tau_{p}(\mathcal{F},B)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,xh,x)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,x)=\mu_{p}(\mathcal{F},B).$
+Moreover by (14) we have the equality $GSV_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
+Thus, if $\nu_{p}(B)=1$ then $\nu_{p}(\tilde{B})=1$, $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)=GSV_{p}(\mathcal{F},B)$ and $\mu_{p}(\tilde{\mathcal{F}},\tilde{B})=\tau_{p}(\tilde{\mathcal{F}},\tilde{B})=GSV_{p}(\tilde{\mathcal{F}},\tilde{B})$, so (11) becomes an equality.
+Moreover, if (11) is an equality, then $B$ is necessarily smooth.
+Indeed, if $B$ were singular, by (10) the number $-\nu_{p}(B)(\nu_{p}(B)-1)+4\alpha$ is positive and from the equality (17) we have that (11) is a strict inequality that is a contradiction.
+Hence the lemma follows.
+∎
 Remark 3.2.
-If $\mathcal{F}$ is a reduced foliation at $p\in\mathbb{C}^{2}$, and $B$ is a smooth separatrix through $p$, then, by Remark 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F}.B)=\tau_{p}(\mathcal{F},B)$.
-Therefore, according to [3, p.
-538], we have $\tau_{p}(\mathcal{F},B)=1$, if $p$ is a non-degenerate singularity, or if $p$ is a saddle-node singularity and $B$ is the strong separatrix; and $\tau_{p}(\mathcal{F},B)=k+1$, if $p$ is a saddle-node singularity and $B$ is the weak separatrix.
-Here $k\in\mathbb{Z}^{+}$ is a formal invariant in the case that $p$ is a saddle-node singularity of $\mathcal{F}$.
-In any cases, $\tau_{p}(\mathcal{F},B)>0$.
+By definition, for any branch $B$ we have $\tau_{p}(B)\leq\mu_{p}(B)$, so after [7, Proposition 6.2] and [7, Propositon 5.7] we get $\tau_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)\leq\mu_{p}(\mathcal{F},B)-GSV_{p}(\mathcal{F},B)$ and
+(19) $\tau_{p}(\mathcal{F},B)\leq\mu(\mathcal{F},B).$
+In addition, by (18), if $B$ is smooth then $\tau_{p}(\mathcal{F},B)=\mu(\mathcal{F},B)$.
 We will now prove the main theorem of this paper.
 3.1. Proof of Theorem A
-Let $\mathcal{F}$ be a holomorphic foliation on $(\mathbb{C}^{2},p)$, and let $B$ be a separatrix of $\mathcal{F}$ at $p$.
-Then, on a minimal reduction of singularities $\pi:(\tilde{X},\mathcal{D})\to(\mathbb{C}^{2},p)$ of $\mathcal{F}$ we marked the reduction process of the pair $(\mathcal{F},B)$:
-(27) $(\mathcal{F}^{(N)},B^{(N)})\overset{\sigma_{N}}{\longrightarrow}\cdots\overset{\sigma_{3}}{\longrightarrow}(\mathcal{F}^{(2)},B^{(2)})\overset{\sigma_{2}}{\longrightarrow}(\mathcal{F}^{(1)},B^{(1)})\overset{\sigma_{1}}{\longrightarrow}(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
-where $\pi=\sigma_{1}\circ\ldots\circ\sigma_{N}$, $B^{(j+1)}$ denotes the strict transform of $B^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $B^{(N)}$ coincide with the normalization of $B$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in B^{(N)}\cap\mathcal{D}$.
-Put $\mu^{(i)}=\mu_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})$, for any $i\in\{0,\ldots,N\}$.
-Applying Lemma 4.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\overset{\sigma_{i+1}}{\longrightarrow}(\mathcal{F}^{(i)},B^{(i)})$, we get
-$3\mu^{(i)}-4\tau^{(i)}=3\mu^{(i+1)}-4\tau^{(i+1)}+\left[\nu_{q^{(i)}}(B^{(i)})(2\nu_{q^{(i)}}(B^{(i)})-m_{q^{(i)}}(\mathcal{F}^{(i)})-1)-4\alpha_{i}\right]$
+Suppose that $\mathcal{F}$ is a holomorphic foliation at $p\in\mathbb{C}^{2}$, $B$ is a separatrix of $\mathcal{F}$ at $p$.
+First, assume that $B$ is singular, that is, $\nu_{p}(B)\geq 2$.
+Then, there is a sequence of blow-ups of $\mathcal{F}$ at $p$ (cf.
+Seidenberg [16]):
+(20) $(\mathcal{F}^{(N)},B^{(N)})\longrightarrow\cdots\longrightarrow(\mathcal{F}^{(2)},B^{(2)})\longrightarrow(\mathcal{F}^{(1)},B^{(1)})\longrightarrow(\mathcal{F}^{(0)},B^{(0)})=(\mathcal{F},B),$
+where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf.
+[11, p.
+1423]).
+Put $\mu^{(i)}:=\mu_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ and $\tau^{(i)}:=\tau_{{q^{(i)}}}(\mathcal{F}^{(i)},B^{(i)})$ for any $i\in\{1,\ldots,N\}$.
+Applying Lemma 3.1 to each blow-up $(\mathcal{F}^{(i+1)},B^{(i+1)})\longrightarrow(\mathcal{F}^{(i)},B^{(i)})$, we get
+$3\mu^{(i)}-4\tau^{(i)}+GSV_{q^{(i)}}(\mathcal{F}^{(i)},B^{(i)})\leq 3\mu^{(i+1)}-4\tau^{(i+1)}+GSV_{q^{(i+1)}}(\mathcal{F}^{(i+1)},B^{(i+1)}).$
 In particular
-$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)$ $=$ $3\mu^{(N)}-4\tau^{(N)}$
-$+$ $\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
-$=$ $3\mu^{(N)}-4\tau^{(N)}+R_{p}(\mathcal{F},B),$
-where the last equality follows from (15).
-We have two possibilities: either $B\in\text{ISO}_{p}(\mathcal{F},B)$ or $B\in\text{Dic}_{p}(\mathcal{F})$.
-If $B\in\text{ISO}_{p}(\mathcal{F},B)$, there exists a component $D\subset\mathcal{D}$ with $D$ non-dicritical such that $\mathcal{F}^{(N)}$ has a reduced singularity at $q^{(N)}\in\mathcal{D}\cap B^{(N)}$.
-In this case, we have $\mu^{(N)}=\tau^{(N)}$ by Remark 4.2, and thus
-$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=-\tau^{(N)}+R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
-If $B\in\text{Dic}_{p}(\mathcal{F})$, then there exists a component $D\subset\mathcal{D}$ with $D$ dicritical such that $B^{(N)}$ is transversal to $D$ at $q^{(N)}$, and $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$.
-In this case, we get
-$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)=R_{p}(\mathcal{F},B)=\text{DG}_{p}(\mathcal{F},B).$
-In both cases, the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ implies $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0 and the theorem follows. $\Box$
-The following example illustrates that the hypothesis $\text{DG}_{p}(\mathcal{F},B)<0$ in Theorem A cannot be removed.
+$3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<3\mu^{(N)}-4\tau^{(N)}+GSV_{q^{(N)}}(\mathcal{F}^{(N)},B^{(N)})=0,$
+where the last equality holds, since $B^{(N)}$ is a smooth curve.
+If $B$ is a smooth curve then, by Remarks 3.2, and 2.1, $GSV_{p}(\mathcal{F},B)=\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$, then the inequality (4) is an equality.
+Reciprocally, if (4) is an equality, then $B$ is necessarily smooth.
+Indeed, if $B$ were singular, by Lemma 3.1, $03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)<$ 0, this is a contradiction.
+Hence, the theorem follows.
+$\Box$
+4. Examples
+In this section, we provide two examples of holomorphic foliations and separatrices that illustrate Theorem A.
+Example 4.1.
+Let $\mathcal{G}_{m,n}$ be the germ of holomorphic foliation at $(\mathbb{C}^{2},0)$ defined by
+$\eta_{m,n}=mxdy-nydx,$
+where $1\leq m\leq n$ are integers. The curve $B:y^{m}-x^{n}=0$ is a separatrix of $\mathcal{G}_{m,n}$ at $0\in\mathbb{C}^{2}$ and
+$GSV_{0}(\mathcal{G}_{m,n},B)$ $=$ $m+n-mn,$
+$\mu_{0}(\mathcal{G}_{m,n},B)$ $=$ $1,$
+$\tau_{0}(\mathcal{G}_{m,n},B)$ $=$ $1$
+We get
+$3\mu_{0}(\mathcal{G}_{m,n},B)-4\tau_{0}(\mathcal{G}_{m,n},B)+GSV_{0}(\mathcal{G}_{m,n},B)=m+n-mn-1=-(m-1)(n-1)\leq 0$
+Thus, Theorem A is verified.
 Example 4.2.
@@
 $\omega_{m}=\left((2m+1)yx^{m+1}+my^{m+2}\right)dx+\left((1-m)xy^{m+1}-2mx^{m+2}\right)dy,$
-where $m\in\mathbb{Z}_{>0}$. Let $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$. We have that $B$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$ such that
+where $m\in\mathbb{Z}_{>0}$. The curve $B:x^{2m+1}+x^{m}y^{m+1}+y^{2m}=0$ is a separatrix of $\mathcal{F}$ at $0\in\mathbb{C}^{2}$. We get
 $GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}3&\text{if $m=1$}\\ -2m^{2}+4m+1&\text{ if $m>1$},\end{array}\right.$
@@
 $\tau_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}5&\text{ if $m=1$}\\ m^{2}+4m+1&\text{if $m>1$}.\end{array}\right.$
-A straightforward calculation reveals that $\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}<\frac{4}{3}$ for $60<m<$ 6, and
-$\lim_{m\to+\infty}\frac{\mu_{0}(\mathcal{F}_{m},B)}{\tau_{0}(\mathcal{F}_{m},B)}=2.$
-On the other hand, we compute $\text{DG}_{0}(\mathcal{F}_{m},B)$ for $m\geq 1$. Note that $\nu_{0}(\mathcal{F}_{m})=m+2$, and $\nu_{0}(B)=2m$, and using [19, equation 4] we get
-$\alpha_{0}=\left\{\begin{array}[]{cl}1&\text{ if $m=1$}\\ m^{2}+m&\text{ if $m>1$}\end{array}\right.$
-After a blow-up at $0\in\mathbb{C}^{2}$, the strict transform of $B$ is a smooth curve, and $m_{0}(\mathcal{F}_{m})=m+2$.
-The reduction of singularities of the pair $(\mathcal{F}_{m},B)$ is achieved in $2m$ steps, and $B\in\text{Dic}_{0}(\mathcal{F}_{m})$.
-Thus
-$\text{DG}_{0}(\mathcal{F}_{1},B)$ $=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F})-1)-4\alpha_{0}\right]$
-$+\left[\nu_{q^{(1)}}(B^{(1)})(2\nu_{q^{(1)}}(B^{(1)})-m_{q^{(1)}}(\mathcal{F}^{(1)})-1)-4\alpha_{1}\right]$
-$=$ $\left[2(4-3-1)-4(1)\right]+\left[1(2-2-1)-4(0)\right]$
-$=$ $-5$
-and for $m>1$, we have
-$\text{DG}_{0}(\mathcal{F}_{m},B)$ $=$ $\sum_{j=0}^{2m-1}\left[\nu_{q^{(j)}}(B^{(j)})(2\nu_{q^{(j)}}(B^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)}_{m})-1)-4\alpha_{j}\right]$
-$=$ $\left[\nu_{0}(B)(2\nu_{0}(B)-m_{0}(\mathcal{F}_{2m})-1)-4\alpha_{0}\right]$
-$+\left[\nu_{q^{(2m-1)}}(B^{(2m-1)})(2\nu_{q^{(2m-1)}}(B^{(2m-1)})-m_{q^{(2m-1)}}(\mathcal{F}^{(2m-1)}_{2m})-1)-4\alpha_{2m-1}\right]$
-$=$ $\left[2m\left(2(2m)-(m+2)-1\right)-4(m^{2}+m)\right]+\left[1(2-2-1)-4(0)\right]$
-$=$ $1\displaystyle 2m^{2}-10m-$ 1 .
-Therefore, $\text{DG}_{0}(\mathcal{F}_{m},B)\geq 0$ if and only if $m\geq 6$. This shows that the hypothesis of Theorem A is optimal.
+A straightforward calculation reveals that
+$3\mu_{0}(\mathcal{F}_{m},B)-4\tau_{0}(\mathcal{F}_{m},B)+GSV_{0}(\mathcal{F}_{m},B)=\left\{\begin{array}[]{cl}-2&\text{if $m=1$}\\ -6m&\text{ if $m>1$}.\end{array}\right.$
+Therefore, Theorem A is verified.
 5. Applications
-In this section, from the proof of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [11] and [19].
+From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17]:
 Corollary 5.1.
-Let $f(x,y)\in\mathbb{C}\{x,y\}$ an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
+Let $f(x,y)\in\mathbb{C}\{x,y\}$ be an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
 $\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
 Proof.
-Consider the singular holomorphic foliation $\mathcal{F}:\omega=df$. In this situation, $C\in\text{ISO}_{0}(\mathcal{F})$. To continue, we compute
-$\text{DG}_{0}(\mathcal{F},C)=R_{0}(\mathcal{F},C)-\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)}),$
-where
-$R_{0}(\mathcal{F},C):=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(2\nu_{q^{(j)}}(C^{(j)})-m_{q^{(j)}}(\mathcal{F}^{(j)})-1)-4\alpha_{j}\right],$
-and $C^{(j+1)}$ denotes the strict transform of $C^{(j)}$ thought $q^{(j)}$ for each $j=0,\ldots,N-1$, $C^{(N)}$ is the normalization of $C$ (or a strict transform of it), $\mathcal{F}^{(i+1)}$ denotes the strict transform of $\mathcal{F}^{(j)}$, and $\mathcal{F}^{(N)}$ is a foliation with a reduced singularity at $q^{(N)}\in C^{(N)}$ as in the proof of Theorem A.
-Since $C^{(N)}$ is smooth, we have $\tau_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})=GSV_{q^{(N)}}(\mathcal{F}^{(N)},C^{(N)})$ by Remark 4.2, and from Lemma 3.1, we deduce
-$\text{DG}_{0}(\mathcal{F},C)=-GSV_{0}(\mathcal{F},C)+\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right]$
-Now, since $\mathcal{F}$ is a curve generalized non-dicritical foliation at $0\in\mathbb{C}^{2}$, and $\text{Sep}_{0}(\mathcal{F})=C$, we have $GSV_{0}(\mathcal{F},C)=0$ by [3, p.
-528].
-Moreover $\mu_{0}(\mathcal{F},C)=\mu(C)$ and $\tau_{0}(\mathcal{F},C)=\tau(C)$ by Remark 2.1.
-Hence
-$\text{DG}_{0}(\mathcal{F},C)=\sum_{j=0}^{N-1}\left[\nu_{q^{(j)}}(C^{(j)})(\nu_{q^{(j)}}(C^{(j)})-1)-4\alpha_{j}\right].$
-Finally, applying successively [19, Claim 4.4] in each step of the resolution of singularities of $C$, and given that the corresponding value $\mathcal{D}_{min}$ of [19, Claim 4.4] in the $j$-th step of the resolution of $C$ is less or equal to $\alpha_{j}$, we get $\text{DG}_{0}(\mathcal{F},C)<0$ implying
-$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}$
-by Theorem A. ∎
+It is a consequence of Remark 2.1 and applying Theorem A to the foliation $\mathcal{F}:df=0$, because $\mu_{p}(\mathcal{F},C)=\mu(C)$, $\tau_{p}(\mathcal{F},C)=\tau(C)$, and $GSV_{p}(\mathcal{F},C)=0$.
+∎
 Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
 Corollary 5.2.
-Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$.
-Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$.
-If $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$ with $\text{DG}_{p}(\mathcal{F},B_{j})<0$ for all $1\leq j\leq\ell$, then
-(28) $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}<\frac{4}{3}.$
+Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
+(21) $03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0
 Proof.
 Suppose that $\mathcal{B}=\sum_{j=1}^{\ell}B_{j}$, where each $B_{j}$ is a separatrix for $\mathcal{F}$. It follows from Theorem A that
-$\sum_{j=1}^{\ell}\left(3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})\right)\leq-\ell<-\ell+1.$
+$\sum_{j=1}^{\ell}\left[3\mu_{p}(\mathcal{F},B_{j})-4\tau_{p}(\mathcal{F},B_{j})+GSV_{p}(\mathcal{F},B_{j})\right]\leq-\ell.$
 Thus
-$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})<-\ell+1=3\ell-4\ell+(4-3)$
-and
-$3\left(\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-\ell+1\right)<4\left(\sum_{j=1}^{\ell}\tau_{p}(\mathcal{F},B_{j})-\ell+1\right),$
+$3\sum_{j=1}^{\ell}\mu_{p}(\mathcal{F},B_{j})-4T_{p}(\mathcal{F},\mathcal{B})+\sum_{j=1}^{\ell}GSV_{p}(\mathcal{F},B_{j})\leq-\ell$
+and, using (7) together with the GSV-index addition formula ([3, p. 29]), we obtain
+$3\left(\mu_{p}(\mathcal{F},\mathcal{B})+\ell-1\right)-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})+2\sum_{i\neq k}i_{p}(B_{i},B_{k})\leq-\ell,$
 so
-$\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
-Since $\ell\geq 1$, we obtain $\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})}\leq\frac{\mu_{p}(\mathcal{F},\mathcal{B})}{T_{p}(\mathcal{F},\mathcal{B})-\ell+1}<\frac{4}{3}.$
+$3\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})\leq-4\ell+3-2\sum_{i\neq k}i_{p}(B_{i},B_{k}).$
+Since $\ell\geq 1$, we get
+$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0 .
 The following example illustrates Corollary 5.2.
@@
 Let $C:x=0$, and consider the effective primitive balanced divisor $\mathcal{B}=C+C_{1,1}$ of separatrices for $\mathcal{F}$ at $0\in\mathbb{C}^{2}$.
-We have $GSV_{0}(\mathcal{F},C)=5$, $GSV_{0}(\mathcal{F},C_{1,1})=5$, $\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$, hence $\frac{\mu_{0}(\mathcal{F},\mathcal{B})}{T_{0}(\mathcal{F},\mathcal{B})}=\frac{25}{26}<\frac{4}{3}$.
+We have $GSV_{0}(\mathcal{F},C)=GSV_{0}(\mathcal{F},C_{1,1})=5$, so
+$GSV_{0}(\mathcal{F},\mathcal{B})=GSV_{0}(\mathcal{F},C)+GSV_{0}(\mathcal{F},C_{1,1})-2i_{0}(C,C_{1,1})=5+5-2\cdot 5=0,$
+$\mu_{0}(\mathcal{F},\mathcal{B})=\mu_{0}(\mathcal{F},C)+\mu_{0}(\mathcal{F},C_{1,1})-deg(\mathcal{B})+1=5+21-2+1=25$, and $T_{0}(\mathcal{F},\mathcal{B})=\tau_{0}(\mathcal{F},C)+\tau_{0}(\mathcal{F},C_{1,1})=5+21=26$.
+Hence
+$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})=3\cdot 25-4\cdot 26=-29<$ 0 .
 We conclude the paper with a question:
-Question 1: Is the inequality (28) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?
+Question: Is the inequality (21) true if instead of $T_{p}(\mathcal{F},\mathcal{B})$ we write $\tau_{p}(\mathcal{F},\mathcal{B})$?
+Acknowledgments. The authors gratefully acknowledge the support of Universidad de La Laguna (Tenerife, Spain), where part of this work was done.
+Funding.
+This work was funded by the Dirección de Fomento de la Investigación at the PUCP through grant DFI-2023-PI0979 and by the Spanish grant PID2019-105896GB-I00 funded by MCIN/AEI/10.13039/501100011033.
+Conflict of interest. The authors of this work declare that they have no conflicts of interest.

### v4

@@
-A Dimca-Greuel type inequality for foliations
+An upper bound for the GSV-index of a foliation
 Abstract.
-Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and $B$ be a separatrix of $\mathcal{F}$.
-We prove the following Dimca-Greuel type inequality $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$, where $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$, $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$, and $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$.
-As a consequence, we provide a new proof of the $\frac{4}{3}$-Dimca-Greuel conjecture for singularities of irreducible plane curve germs, with foliations ingredients, that differs from those given by Alberich-Carramiñana, Almirón, Blanco, Melle-Hernández and Genzmer-Hernandes, but it is in line with the idea developed by Wang.
+Let $\mathcal{F}$ be a holomorphic foliation at $p\in\mathbb{C}^{2}$, and let $B$ be a separatrix of $\mathcal{F}$.
+We prove the following upper bound $GSV_{p}(\mathcal{F},B)\leq 4\tau_{p}(\mathcal{F},B)-3\mu_{p}(\mathcal{F},B)$, where $GSV_{p}(\mathcal{F},B)$ is the Gómez-Mont-Seade-Verjovsky index of the foliation $\mathcal{F}$ with respect to $B$, $\mu_{p}(\mathcal{F},B)$ is the multiplicity of $\mathcal{F}$ along $B$ and $\tau_{p}(\mathcal{F},B)$ is the dimension of the quotient of $\mathbb{C}\{x,y\}$ by the ideal generated by the components of any $1$-form defining $\mathcal{F}$ and any equation of $B$.
 Key words and phrases:
-Holomorphic foliations, Milnor number, Tjurina number, Multiplicity of a foliation along a divisor of separatrices, Dimca-Greuel bound
+Holomorphic foliations, Gómez-Mont-Seade-Verjovsky index, Tjurina number, Multiplicity of a foliation along a divisor of separatrices.
 2020 Mathematics Subject Classification:
@@
 1. Introduction
-Let $\mathbb{C}\{x,y\}$ be the ring of complex convergent power series in two variables.
-Consider the germ of isolated plane curve singularity $C$ defined by the non-unit $f(x,y)\in\mathbb{C}\{x,y\}$.
-The Milnor number of $C$ is
-$\mu(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(\partial_{x}f,\partial_{y}f)$
-and its Tjurina number is
-$\tau(C)=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(f,\partial_{x}f,\partial_{y}f).$
-Is clear from their definitions that $\tau(C)\leq\mu(C)$.
-In [6, Question 4.2], Dimca and Greuel present a family of curves $\{C_{m}\}_{m\geq 1}$ such that the quotient $\frac{\mu(C_{m})}{\tau(C_{m})}$ is strictly increasing with limit $4/3$ and they ask if this property is true for any plane curve singularity.
-The question was affirmatively answered in [1], [10], and [17] for the irreducible case and in [2] for the reduced case.
-It is natural to wonder whether this bound also holds if we consider germs of singular foliations instead of plane curve germs.
-More precisely, if $\mathcal{F}$ is a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, with $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, then the Milnor number of $\mathcal{F}$ is by definition the dimension of $\mathbb{C}\{x,y\}/(P,Q)$ and we denote it by $\mu_{p}(\mathcal{F})$.
-Moreover, if $\mathcal{B}=\mathcal{B}_{0}-\mathcal{B}_{\infty}$ is a balanced divisor of separatrices of $\mathcal{F}$ (see [9, p.
-4], [11, Definition 3.1]), the Tjurina number of $\mathcal{F}$ with respect to $\mathcal{B}_{0}:f(x,y)=0$ is the dimension
-$\tau_{p}(\mathcal{F},\mathcal{B}_{0}):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
-If $\mathcal{F}$ is the hamiltonian foliation given by the $1$-form $df$ for some $f\in\mathbb{C}\{x,y\}$, then the ratio
-(1) $\frac{\mu_{p}(\mathcal{F})}{\tau_{p}(\mathcal{F},\mathcal{B}_{0})}$
-equals $\frac{\mu_{p}(C)}{\tau_{p}(C)}$, where $C$ is the germ of the curve of equation $f(x,y)=0$.
-If the foliation $\mathcal{F}$ is of the second type (see [9, Definition 2.5]) and $\mathcal{B}$ is a balanced divisor of separatrices for $\mathcal{F}$ then, in [8, Proposition 4.1], we have proved the inequality $\mu_{p}(\mathcal{F})\leq 2\tau_{p}(\mathcal{F},\mathcal{B}_{0})$  generalizing to foliations of second type this well-known inequality in the context of germs of reduced plane curves (see for instance [14] where the general case of hypersurfaces is studied).
-It is natural to wonder if we can improve the bound $2$ for the ratio (1).
-We note that, in general, the Dimca-Greuel bound does not hold for foliations, as the following example illustrates: if we borrow from [7, Example 6.5] the family of dicritical foliations $\mathcal{F}_{k}$, $k\geq 3$, $\0\lambda\in\mathbb{C}\backslash\{0\}$, given by the $1$-form
-$\omega_{k}=y\left(2x^{2k-2}+2(\lambda+1)x^{2}y^{k-2}-y^{k-1}\right)dx+x\left(y^{k-1}-(\lambda+1)x^{2}y^{k-2}-x^{2k-2}\right)dy,$
-then for the effective balanced divisor of separatrices $\mathcal{B}=\mathcal{B}_{0}:xy=0$ of $\mathcal{F}_{k}$ we have $\tau_{0}(\mathcal{F}_{k},\mathcal{B}_{0})=3k-2$ and $\mu_{0}(\mathcal{F}_{k})=k(2k-1)$; hence for the family $\mathcal{F}_{k}$ of foliations the ratio (1) goes to infinity when $k$ grows.
-On the other hand, following Camacho-Lins Neto-Sad [4, Section 4], we can consider the multiplicity of the singular foliation $\mathcal{F}$ along a separatrix $B$ of $\mathcal{F}$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition).
-We note that if $B=\{f(x,y)=0\}$, and $\mathcal{F}:df=0$, then $\mu_{p}(\mathcal{F},B)$ also coincides with the Milnor number $\mu(B)$ of $B$ (see Remark 2.1).
-Therefore, we could think to studying if the inequality
-$03\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)<$ 0
-holds. But, for the germ of foliation $\mathcal{F}$ at $(\mathbb{C}^{2},0)$ defined by
-(2) $\omega=y\left(13x^{7}+6y^{7}\right)dx-x\left(5y^{7}+12x^{7}\right)dy,$
-and the separatrix
-(3) $B:x^{13}+x^{6}y^{7}+y^{12}=0,$
-we have $\mu_{0}(\mathcal{F},B)=85$ and $\tau_{0}(\mathcal{F},B)=61$. Thus
-$113\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61=255-244=$ 11 .
-In order to establish an inequality of the Dimca-Greuel’s type for foliations, it is necessary to add the Gómez-Mont–Seade–Verjovsky index (in short GSV-index) (see [12]) as stating our main result.
+Let $\mathcal{F}$ be a germ of singular holomorphic foliation at $(\mathbb{C}^{2},p)$ given by the $1$-form $\omega:=P(x,y)dx+Q(x,y)dy$, where $P(x,y),Q(x,y)\in\mathbb{C}\{x,y\}$, and $B$ be a separatrix of $\mathcal{F}$.
+Several numerical invariants can be studied for the pair $(\mathcal{F},B)$, such as the Gómez-Mont–Seade–Verjovsky index (see [10]) of $\mathcal{F}$ with respect to $B$, denoted by $GSV_{p}(\mathcal{F},B)$, the multiplicity of $\mathcal{F}$ along $B$, $\mu_{p}(\mathcal{F},B)$ (see Section 2 for definition), and the Tjurina number of $\mathcal{F}$ with respect to $B:f(x,y)=0$, defined by
+$\tau_{p}(\mathcal{F},B):=\dim_{\mathbb{C}}\mathbb{C}\{x,y\}/(P,Q,f).$
+In this paper, inspired by the blow-up techniques developed by Wang [14], we establish an optimal upper bound for $GSV_{P}(\mathcal{F},B)$ in terms of $\mu_{p}(\mathcal{F},B)$ and $\tau_{p}(\mathcal{F},B)$ as follows:
 Theorem A.
 Let $\mathcal{F}$ be a germ of a singular holomorphic foliation on $(\mathbb{C}^{2},p)$, and $B$ be a separatrix of $\mathcal{F}$ at $p$. Then
-(4) $3\mu_{p}(\mathcal{F},B)-4\tau_{p}(\mathcal{F},B)+GSV_{p}(\mathcal{F},B)\leq 0$
+(1) $GSV_{p}(\mathcal{F},B)\leq 4\tau_{p}(\mathcal{F},B)-3\mu_{p}(\mathcal{F},B)$
 where $GSV_{p}(\mathcal{F},B)$ denotes the GSV-index of $\mathcal{F}$ with respect to $B$. Moreover, the equality holds if and only if $B$ is smooth.
-We will prove Theorem A in Section 3. For the foliation given in (2), and $B$ as in (3), we have $GSV_{0}(\mathcal{F},B)=-47$. Hence,
-$03\mu_{0}(\mathcal{F},B)-4\tau_{0}(\mathcal{F},B)+GSV_{0}(\mathcal{F},B)=3\cdot 85-4\cdot 61+(-47)=-36<$ 0
-verifying Theorem A.
-As a consequence of Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17] (see Corollary 5.1).
+Note that if $B=\{f=0\}$ is an irreducible plane curve germ, then by applying Theorem A to the foliation $\mathcal{F}:df=0$, we obtain the bound stated in [6, Question 4.2] and proved simultaneously in [1], [8] and [14], since $\mu_{p}(\mathcal{F},B)=\mu(B)$, $\tau_{p}(\mathcal{F},B)=\tau(B)$, and $GSV_{p}(\mathcal{F},B)=0$ in this case.
+Here, $\mu(B)$ and $\tau(B)$ denote the classical Milnor and Tjurina numbers of $B$.
+A complete answer to this question was given by Almirón [2].
+Moreover, the author proposed a broader perspective of the study of the difference between Milnor and Tjurina numbers for isolated complete intersection singularities.
 2. Multiplicity of a foliation along a divisor of separatrices
@@
 By convention we put $\mu_{p}(\mathcal{F},\mathcal{B})=1$ for any empty divisor $\mathcal{B}$.
-In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [13, Equation (2.2), p.
+In particular, when $\mathcal{B}=B_{1}+\cdots+B_{r}$ is an effective divisor of separatrices of the singular foliation $\mathcal{F}$ at $p$, we rediscover [11, Equation (2.2), p.
 329] (for the reduced plane curve $C:=\cup_{i=1}^{r}B_{i}$).
@@
 (5) $\mu_{p}(\mathcal{F},B)-\tau_{p}(\mathcal{F},B)=\mu_{p}(B)-\tau_{p}(B).$
-Consequently when $B$ is quasihomogeneous by ([15], Satz p.123 a) and d)] and (8) or by [18, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
+Consequently when $B$ is quasihomogeneous by ([12], Satz p.123 a) and d)] and (5) or by [15, Theorem 4], $B$ is analytically equivalent to $y^{n}+x^{m}=0$ for some natural coprime numbers $n$, $m>1$, in this case, we get $\mu_{p}(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)$.
 3. Proof of Theorem A
@@
 If $C$ is singular then $G$ is not zero and $\bar{\phi}_{*}G$ is a finite $\mathcal{O}_{T}$-module and after Wang ([14]) we put $\mathcal{D}(t)=\hbox{\rm length}_{\mathcal{O}_{T,t}}((\bar{\phi}_{*}G)_{t})$ and $\alpha:=\min_{t\in T}\mathcal{D}(t)$.
-By [17, Claim 4.4] we have
+By [14, Claim 4.4] we have
 (7) $\alpha>\dfrac{\nu_{p}(C)(\nu_{p}(C)-1)}{4},\,\,\,\,\text{for}\,\,\,\,\nu_{p}(C)\geq 2.$
@@
 (11) $GSV(\mathcal{F},B)=\tau_{p}(\mathcal{F},B)-\tau_{p}(B).$
-From [17, Equation (4)] we obtain
+From [14, Equation (4)] we obtain
 (12) $\tau_{p}(B)-\tau_{q}(\tilde{B})\geq\frac{\nu_{p}(B)(\nu_{p}(B)-1)}{2}+\alpha,$
@@
 where $B^{(i+1)}$ denotes the strict transform of $B^{(i)}$ thought $q^{(i)}$ for each $i=0,\ldots,N-1$, $B^{(N)}$ is smooth, and $\mathcal{F}^{(N)}$ is a foliation with a simple singularity at $q^{(N)}\in B^{(N)}$, or $\mathcal{F}^{(N)}$ is a regular foliation at $q^{(N)}$, and $B^{(N)}$ is a dicritical separatrix (cf.
-[11, p.
+[9, p.
 1423]).
@@
 5. Applications
-From Theorem A we deduce the Dimca-Greuel bound for the case of irreducible plane curve germs, proved simultaneously in [1], [10] and [17]:
-Corollary 5.1.
-Let $f(x,y)\in\mathbb{C}\{x,y\}$ be an irreducible power series with an isolated singularity at $0\in\mathbb{C}^{2}$. Put $C:f(x,y)=0$ the branch defined by $f$. Then
-$\frac{\mu(C)}{\tau(C)}<\frac{4}{3}.$
-Proof.
-It is a consequence of Remark 2.1 and applying Theorem A to the foliation $\mathcal{F}:df=0$, because $\mu_{p}(\mathcal{F},C)=\mu(C)$, $\tau_{p}(\mathcal{F},C)=\tau(C)$, and $GSV_{p}(\mathcal{F},C)=0$.
-∎
-Now, if in (1), we change $\mu_{p}(\mathcal{F})$ to the multiplicity of the singular foliation $\mathcal{F}$ along a reduced curve $C$, $\mu_{p}(\mathcal{F},C)$ (see Section 2), and we put in the denominator of (1) the sum $T_{p}(\mathcal{F},C):=\sum_{B}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality of the Dimca-Greuel type for foliations:
+We define the sum $T_{p}(\mathcal{F},C):=\sum_{B\subset C}\tau_{p}(\mathcal{F},B)$ of Tjurina numbers of $\mathcal{F}$ along the irreducible components of $C$, then we get the following inequality for foliations:
 Corollary 5.1.
 Let $\mathcal{F}$ be a germ of a singular holomorphic foliation at $(\mathbb{C}^{2},p)$. Let $\mathcal{B}$ be an effective primitive balanced divisor of separatrices for $\mathcal{F}$ at $p$. Then
-(21) $03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})<$ 0
+(18) $GSV_{p}(\mathcal{F},\mathcal{B})<4T_{p}(\mathcal{F},\mathcal{B})-3\mu_{p}(\mathcal{F},\mathcal{B})$
 Proof.
@@
 Hence
-$03\mu_{p}(\mathcal{F},\mathcal{B})-4T_{p}(\mathcal{F},\mathcal{B})+GSV_{p}(\mathcal{F},\mathcal{B})=3\cdot 25-4\cdot 26=-29<$ 0 .
+$GSV_{p}(\mathcal{F},\mathcal{B})=0<4T_{p}(\mathcal{F},\mathcal{B})-3\mu_{p}(\mathcal{F},\mathcal{B})=29.$
 We conclude the paper with a question: